### v3

# Rational powers, invariant ideals, and the summation formula

Abstract.
We provide explicit descriptions for the rational powers and Rees valuations of several classes of ideals invariant under natural actions of tori and products of general linear groups, in terms of polyhedra and lattice points. This allows us to show that a version of Mustaţă-Takagi’s summation formula for multiplier ideals also holds for the rational powers of these ideals. Moreover, for arbitrary ideals in normal domains that are finitely generated over algebraically closed fields, we prove a weaker version of this formula that holds for sufficiently large rational numbers.
Key words and phrases:
Rational powers, integral closures, Rees valuations, multiplier ideals, invariant ideals.
2020 Mathematics Subject Classification:
Primary: 13B22, 13A50, 13A18. Secondary: 14E05, 05E40.
1 The third author was partially supported by the Simons Foundation.
2 The fourth author was partially funded by NSF Grant DMS #2001645/2303605 and DMS #2401522.
1. Introduction
Rees valuations and rational powers of ideals play a central role in many results in algebraic geometry and commutative algebra [36, 39, 20, 13, 12, 46, 27, 43, 3, 26, 52]. These two sets of invariants are so closely related that each uniquely determines the other (see Section missing 2). However, explicitly computing them is a difficult task because it requires determining the order valuations for every prime divisor arising in the exceptional divisor of the normalized blowup of the ideal [50]. In fact, complete descriptions of Rees valuations in the literature are limited to monomial ideals, where they correspond to supporting hyperplanes of Newton polyhedra [38, Theorem 10.3.5]. One of the primary goals of this paper is to provide explicit descriptions of both Rees valuations and rational powers for several classes of ideals in polynomial rings that are invariant under natural actions of tori and products of general linear groups.
Originally considered by Samuel and Rees in the 1950s [54, 50], the notion of rational powers of an ideal $I$ in a Noetherian domain $R$ was introduced by Lejeune-Jalabert and Teissier [45, §4.2]. The power $\overline{I^{w}}$ for $w\in\mathbb{Q}_{\geqslant 0}$ is defined as $\overline{I^{w}}:=\cap_{v}\{x\mid v(x)\geqslant wv(I)\}$, where the intersection is taken over all the discrete valuations of $\operatorname{Quot}(R)$ that are nonnegative in $R$, or alternatively, over the Rees valuations of $I$. This construction provides a powerful tool for understanding the structure of filtrations of ideals such as integral closures [36, 39], symbolic powers [26], Frobenius powers [52], and multiplier ideals [44, §9.6]. Despite its historical roots, rational powers have seen limited development, as noted by Knutson in 2010, who observed that this “beautiful construction has sat essentially unused in the literature for over 50 years now" [42].
Let $X=X_{m\times n}$ be a generic $m\times n$ matrix, meaning its entries are distinct variables. The polynomial ring $\mathbb{C}[X]$ carries a natural action of $\mathrm{GL}_{m}(\mathbb{C})\times\mathrm{GL}_{n}(\mathbb{C})$ defined by $(A,B)\cdot X=A^{-1}XB$, where $A\in\mathrm{GL}_{m}(\mathbb{C})$ and $B\in\mathrm{GL}_{n}(\mathbb{C})$. The ideals in the polynomial ring $\mathbb{C}[X]$ that are invariant under this action have been extensively studied in the literature, see e.g. [17, 18, 6, 10, 35, 40, 21, 7]. The combination of representation theory and the theory of algebras with straightening laws has led to a wide range of explicit results about the structure and properties of these ideals. If one considers a symmetric generic matrix $Y=Y_{n\times n}$ and $\operatorname{GL}_{n}(\mathbb{C})$-action on $\mathbb{C}[Y]$ given by $A\cdot Y=AYA^{t}$, or a skew-symmetric generic matrix $Z=Z_{n\times n}$ with $\operatorname{GL}_{n}(\mathbb{C})$-action on $\mathbb{C}[Z]$ given by $A\cdot Z=AZA^{t}$, one obtains analogous well-studied invariant ideals, see e.g. the references above and [2, 1, 19, 14].
We introduce the concept of Rees package, which is a combinatorial characteristic-free construction that encapsulates the description of both Rees valuations and rational powers in terms of polyhedral data (see Section missing 3). In our first theorem, we show that the classes of invariant ideals described above admit such a Rees package. Furthermore, in arbitrary characteristic, we show that monomial ideals in affine semigroup rings, sums of products of determinantal ideals of generic, symmetric, and skew-symmetric matrices, and products of determinantal ideals of Hankel matrices, also admit a Rees package.
Theorem A.
Let $\mathbb{k}$ be an arbitrary field. The following classes of ideals admit a Rees package. In particular, their rational powers and Rees valuations can be explicitly described in terms of lattice points and supporting hyperplanes of polyhedra.
  1. (a)
    (Theorem 3.1) Monomial ideals in affine semigroup rings.
  2. (b)
    (Theorem 3.2, §missing 3.2) Sums of products of determinantal ideals in $\mathbb{k}[X_{m\times n}]$ of a generic matrix $X_{m\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{C}[X_{m\times n}]$.
  3. (c)
    (Theorem 3.3, §missing 3.3) Sums of products of determinantal ideals in $\mathbb{k}[Y_{n\times n}]$ of a generic symmetric matrix $Y_{n\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{k}[Y_{n\times n}]$
  4. (d)
    (Theorem 3.4, §missing 3.4) Sums of products of ideals of Pfaffians in $\mathbb{k}[Z_{n\times n}]$ of a skew-symmetric generic matrix $Z_{n\times n}$. If $\mathbb{k}=\mathbb{C}$, more generally any $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideal of $\mathbb{C}[Z_{n\times n}]$.
  5. (e)
    (Theorem 3.5) Products of determinantal ideals in $\mathbb{k}[W]$ of a Hankel matrix of variables $W$.
In [49], Mustaţă showed a summation formula for multiplier ideals of pairs ideal sheaves, and this formula was later generalized by Takagi [56] and by Jow-Miller [41]. Mustaţă-Takagi summation formula raises the natural question of whether similar identities hold for symbolic powers and rational powers of ideals. The former was established in [32, 31] and the latter in [4] for monomial ideals. Thus we propose the following conjecture (cf. [4, Question 2.4]).
[…]
Let $\mathbb{k}$ be an algebraically closed field and $R$ and $S$ normal domains finitely generated over $\mathbb{k}$. Set $T:=R\otimes_{\mathbb{k}}S$. The following formula holds for every $w\in\mathbb{Q}_{\geqslant 0}$ and nonzero ideals $I\subset R$ and $J\subset S$
$\overline{\left(IT+JT\right)^{w}}=\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}\,T.$
We devote the last two sections of the paper to tackle this conjecture. First, we note that the restriction on the field is partially justified by a related question (see Section missing 4 and Section missing 4). We also note that the use of two different $\mathbb{k}$-algebras is essential in this conjecture, as the summation formula for rational powers does not hold in general if instead one considers two ideals in the same ring, see Section missing 4.
In Section missing 4 we show that if two ideals $I\subset R$ and $J\subset S$ admit a Rees package, e.g. the invariant ideals in Theorem A, then $IT+JT$ does as well. Moreover, a Rees package for $IT+JT$ can be obtained by ‘gluing’ those of $I$ and $J$. In particular, this shows that the Rees valuations of $IT+JT$ are obtained by pairing the ones of $I$ and $J$ (see Section missing 4). This result allows us to compute rational powers of more classes of ideals, see for instance Section missing 4.
In our next theorem, we use Section missing 4 to settle Section missing 1 for ideals with a Rees package without restrictions on the field.
Theorem B (Theorem 4.1, Section missing 4).
Adopt the notation in Section missing 1 with $\mathbb{k}$ an arbitrary field. The conclusion of Section missing 1 holds if $I$ and $J$ admit a Rees package. In particular, this holds if $I$ and $J$ belong to any of the classes of ideals in Theorem A, and not necessarily both from the same class.
In the following theorem, we provide further evidence for Section missing 1. Here we prove an inclusion for arbitrary ideals and $w\gg 0$. This result can be seen as a weaker version of Section missing 1 since we always have
$\sum_{\stackrel{{\scriptstyle 0\leqslant\tau\leqslant w,}}{{{\scriptscriptstyle\tau\in\mathbb{Q}}}}}\overline{I^{\tau}}\,\overline{J^{w-\tau}}\,T=\bigcap_{\stackrel{{\scriptstyle 0\leqslant\tau\leqslant w,}}{{{\scriptscriptstyle\tau\in\mathbb{Q}}}}}\left(\overline{I^{\tau}}T+\overline{J^{w-\tau}}T\right),$
see Section missing 5.
Theorem C (Theorem 5.1).
Following the notation in Section missing 1, there exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}_{\geqslant 0}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}\,T\subseteq\overline{\left(IT+JT\right)^{w}}\subseteq\overline{I^{\tau}}T+\overline{J^{w-\tau}}T.$
[…]
We now outline the content of the paper. Section missing 2 includes general details about rational powers and Rees valuations, as well as the definition of rational powers of ideals sheaves. Section missing 3 contains the definition of Rees package and the proof of Theorem A. Section missing 4 includes the proof of Theorem B. This section also includes examples where our results are applied. Section missing 5 is devoted to the proof of Theorem C.
Acknowledgments
The authors are grateful to Robert Lazarsfeld for several enlightening insights about multiplier ideals, integral closures, and vanishing theorems. We also thank Claudiu Raicu for helpful correspondence related to the proof of Theorem 3.2 and Bernard Teissier for pointing out §missing 3.1. The fourth author thanks Alessandro De Stefani and Alexandra Seceleanu for discussions related to Section missing 2. The third author was partially supported by the Simons Foundation. The fourth author was partially funded by NSF Grant DMS #2001645/2303605 and DMS #2401522. We acknowledge support by the NSF Grant DMS #1928930, while the second and fourth authors were in residence at the Simons Laufer Mathematical Science Institute (formerly MSRI) in Berkeley, California, during the Spring 2024 semester.
2. Rational powers and Rees valuations
In this section we review and extend some relevant facts about rational powers and Rees valuations. Our main references in this subject is [38]. We note that several of the statements in this section have been previously established for the integral closures of powers of ideals. Our goal is to show that these properties continue to hold in the more general setting of rational powers and to provide references for the corresponding assertions. We restrict our exposition to the following setup where all of our results occur.
[…]
Let $(R,\mathfrak{m},\mathbb{k})$ be a Noetherian local ring, or positively graded over $\mathbb{k}$ with $\mathfrak{m}=R_{>0}:=\oplus_{n\in{\mathbb{Z}}_{>0}}R_{n}$. We also assume $R$ is an integral domain and denote by $\operatorname{Quot}(R)$ its field of fractions. We set $I\subset R$ to be an ideal.
[…]
Adopt Section missing 2. An element $x\in R$ is said to be integral over $I$ if there exists $n\in{\mathbb{Z}}_{\geqslant 0}$ and $a_{i}\in I^{i}$ such that $x^{n}+a_{1}x^{n-1}+a_{2}x^{n-2}+\cdots+a_{n}=0$. The set of all such element is an ideal denoted by $\overline{I}$ and called the integral closure of $I$. Let $w=\frac{p}{q}$ be a nonnegative rational number with $p\in{\mathbb{Z}}_{\geqslant 0}$ and $q\in{\mathbb{Z}}_{>0}$. The $w$-th rational power of $I$ is the ideal defined by
$\overline{I^{w}}:=\{x\in R\mid x^{q}\in\overline{I^{p}}\}.$
[…]
It follows from the definitions that $\overline{I^{w}}$ is well-defined, i.e., if $\frac{p}{q}=\frac{p^{\prime}}{q^{\prime}}$ are two representations of $w$, then $\{x\in R\mid r^{q}\in\overline{I^{p}}\}=\{x\in R\mid r^{q^{\prime}}\in\overline{I^{p^{\prime}}}\}$. Furthermore, $\overline{I^{0}}=R$ and if $n\in{\mathbb{Z}}_{\geqslant 0}$ then $\overline{I^{\frac{n}{1}}}=\overline{I^{n}}$.
In the following proposition we show that rational powers localize.
[…]
Adopt Section missing 2. For every multiplicative subset $W\subset R$ and rational number $w\in{\mathbb{Q}}_{\geqslant 0}$ we have $W^{-1}\overline{I^{w}}=\overline{(W^{-1}I)^{w}}$.
Proof.
If $w$ is an integer, the statement follows from [38, Proposition 10.5.5]. Now let $w$ be arbitrary and write $w=\frac{p}{q}$. If $x\in\overline{I^{w}}$, then $(W^{-1}x)^{q}=W^{-1}x^{q}\in W^{-1}\overline{I^{p}}=\overline{(W^{-1}I)^{p}}$ and so $W^{-1}x\in\overline{(W^{-1}I)^{w}}$. On the other hand, if $y\in\overline{(W^{-1}I)^{w}}$, then $y^{q}\in\overline{(W^{-1}I)^{p}}=W^{-1}\overline{I^{p}}$. Therefore there exists $a\in W$ such that $ay^{q}\in\overline{I^{p}}$. The latter implies $(ay)^{q}\in\overline{I^{p}}$ and so $ay\in\overline{I^{w}}$. Thus $y\in W^{-1}\overline{I^{w}}$, finishing the proof. ∎
A valuation is discrete if it is integer valued. Integral closures and rational powers can also be defined via discrete valuations: by [38, Theorem 6.8.3(3)] we have that $x\in\overline{I}$ if and only if $x\in IV$ for every discrete valuation ring $R\subseteq V\subseteq\operatorname{Quot}(R)$. Moreover, the latter condition can be verified by using finitely many valuations, the Rees valuations of $I$.
[…]
Adopt Section missing 2. Let $I=(a_{1},\dots,a_{d})$ be a set of generators and, for each $i=1,\dots,d$, set $R_{i}=R\left[\frac{a_{1}}{a_{i}},\ldots,\frac{a_{d}}{a_{i}}\right]$ and $\overline{R_{i}}$ its integral closure in $\operatorname{Quot}(R_{i})=\operatorname{Quot}(R)$. Let ${\operatorname{\mathscr{RV}}}(I)$ be the set of distinct discrete valuation rings arising as $V=\left(\overline{R_{i}}\right)_{\mathfrak{p}}$, where $\mathfrak{p}$ varies over the minimal primes of $I\overline{R_{i}}=a_{i}\overline{R_{i}}$. The set ${\operatorname{\mathscr{RV}}}(I)$ is called the Rees valuations rings of $I$. We will often abuse notation and identify each $V\in{\operatorname{\mathscr{RV}}}(I)$ with any valuation $v$ which is a positive integer multiple of the order valuation of $V$.
We say that a discrete valuation $v$ of $\operatorname{Quot}(R)$ is centered in $R$, or simply that $v$ is a valuation of $R$, if $v(x)\geqslant 0$ for every $x\in R$. If $v$ is a valuation of $R$, then set $v(I)=\min\{v(x)\mid x\in I\}$.
[…]
Let ${\operatorname{\mathscr{RV}}}(I)=\{v_{1},\dots,v_{r}\}$ and set $e:={\operatorname{lcm}}(v_{1}(I),\dots,v_{r}(I))$ (or any multiple of it). For every $w\in{\mathbb{Q}}_{\geqslant 0}$ and $n:=\lceil we\rceil$ we have $\overline{I^{w}}=\overline{I^{\frac{n}{e}}}$ (see [38, Proposition 10.5.5]). Moreover, $\overline{I^{\frac{1}{e}}}=\sqrt{I}$ [38, Example 10.5.4].
As mentioned above, Rees valuations can be used to define integral closures.
Theorem 2.1 ([38, Theorem 10.2.2]).
Adopt Section missing 2 and the notation from Section missing 2. For every $n\in{\mathbb{Z}}_{\geqslant 0}$ we have
$\overline{I^{n}}=\cap_{i=1}^{d}\left(I^{n}\overline{R_{i}}\cap R\right)=\cap_{i=1}^{d}\left(a_{i}^{n}\overline{R_{i}}\cap R\right)=\cap_{V\in{\operatorname{\mathscr{RV}}}(I)}\left(I^{n}V\cap R\right)=\{x\in R\mid v(x)\geqslant nv(I),\forall v\in{\operatorname{\mathscr{RV}}}(I)\},$
and no $V$ or $v$ can be omitted without breaking some of these equalities.
The following lemma shows that rational powers can be defined via the same valuations defining integral closures.
[…]
Adopt Section missing 2 and the notation from Section missing 2. Let $\mathscr{T}$ be a set of discrete valuations of $R$. Then $\overline{I^{n}}=\{x\in R\mid v(x)\geqslant nv(I),\forall v\in\mathscr{T}\}$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$ if and only if $\overline{I^{w}}=\{x\in R\mid v(x)\geqslant wv(I),\forall v\in\mathscr{T}\}$ for every $w\in{\mathbb{Q}}_{\geqslant 0}$. In particular, for every $w\in{\mathbb{Q}}_{\geqslant 0}$, we have
$\overline{I^{w}}=\{x\in R\mid v(x)\geqslant wv(I),\forall v\in{\operatorname{\mathscr{RV}}}(I)\},$
and no $v$ can be omitted without breaking some of these equalities.
Proof.
The backwards implication is trivial. For the forward implication, fix $w\in{\mathbb{Q}}_{\geqslant 0}$ and write $w=\frac{p}{q}\in{\mathbb{Q}}_{\geqslant 0}$ with $p\in{\mathbb{Z}}_{\geqslant 0}$ and $q\in{\mathbb{Z}}_{>0}$. Since $x\in\overline{I^{w}}$ if and only if $x^{q}\in\overline{I^{p}}$, and this is the case if and only if $qv(x)\geqslant pv(I)$ for every $v\in\mathscr{T}$, the first statement follows. The second statement follows from the first one and Theorem 2.1. ∎
The previous two results lead to the following useful fact about associated primes of rational powers.
[…]
Adopt Section missing 2. For each $V\in{\operatorname{\mathscr{RV}}}(I)$, set $\mathfrak{p}_{V}:=\{x\in R\mid v(x)>0\}$. We have
$\cup_{w\in{\mathbb{Q}}_{\geqslant 0}}\operatorname{Ass}(R/\overline{I^{w}})=\cup_{n\in{\mathbb{Z}}_{\geqslant 0}}\operatorname{Ass}(R/\overline{I^{n}})=\{\mathfrak{p}_{V}\mid V\in{\operatorname{\mathscr{RV}}}(I)\}.$
Proof.
The second equality is proved in [38, Discussion 10.1.3] and the inclusion $\supseteq$ in the first one is trivial. It remains to show the first inclusion $\subseteq$. For each $V\in{\operatorname{\mathscr{RV}}}(I)$, let $\eta_{V}$ be its maximal ideal. From Section missing 2 we obtain $\overline{I^{w}}=\cap_{V\in{\operatorname{\mathscr{RV}}}(I)}(\eta_{V}^{\lceil wv(I)\rceil}\cap R)$. The result follows as each ideal $\eta_{V}^{\lceil wv(I)\rceil}\cap R$ is $\mathfrak{p}_{V}$-primary. ∎
The following is a useful characterization of Rees valuations.
[…]
Adopt Section missing 2. Let $I\subseteq R$ and $v_{1},\dots,v_{r}$ be discrete valuations of $R$. Then the following statements are equivalent.
  1. (i)
    ${\operatorname{\mathscr{RV}}}(I)=\{v_{1}\ldots,v_{r}\}$.
  2. (ii)
    $\overline{I^{n}}=\{x\in R\mid v_{i}(x)\geqslant nv_{i}(I),\forall i=1,\ldots,r\}$ for all $n\in{\mathbb{Z}}_{\geqslant 0}$, and no $v_{i}$ can be omitted without breaking some of these equalities.
  3. (iii)
    $\overline{I^{w}}=\{x\in R\mid v_{i}(x)\geqslant wv_{i}(I),\forall i=1,\ldots,r\}$ for all $w\in{\mathbb{Q}}_{\geqslant 0}$, and no $v_{i}$ can be omitted without breaking some of these equalities.
Proof.
The implication (i) $\Rightarrow$ (ii) follows from Theorem 2.1, and (ii) $\Rightarrow$ (iii) from Section missing 2. It remains to show (iii) $\Rightarrow$ (i), but this is content of [38, Theorem 10.1.6], finishing the proof. ∎
The next example shows a particular situation where the previous proposition directly applies to compute Rees valuations and rational powers of certain classes of ideals.
[…]
The equivalence (i) $\Leftrightarrow$ (iii) in Section missing 2 shows that if one knows the Rees valuations of an ideal, then one can compute its rational power; and conversely, if one knows the rational powers of the ideal, then one can determine its Rees valuations.
[…]
Let $R$ be a regular ring and $I\subseteq R$ an ideal. Assume that there exist prime ideals $\mathfrak{p}_{1},\ldots,\mathfrak{p}_{r}$ and positive integers $a_{1},\ldots,a_{r}$ such that $\overline{I^{n}}=\cap_{i=1}^{r}\mathfrak{p}_{i}^{(na_{i})}$ is the primary decomposition of $\overline{I^{n}}$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$ that is irredundant for some $n$; here $\mathfrak{p}_{i}^{(m)}$ denotes the $m$th-symbolic power of $\mathfrak{p}_{i}$. Since $R_{\mathfrak{p}_{i}}$ is a regular local ring for every $i$, the functions $v_{i}:R\to{\mathbb{Z}}_{\geqslant 0}$ given by $v_{i}(x)=\max\big{\{}n\mid x\in\mathfrak{p}_{i}^{(n)}\big{\}}$ extend to discrete valuations of $\operatorname{Quot}(R)$. Therefore, by Section missing 2 one has ${\operatorname{\mathscr{RV}}}(I)=\{v_{1},\ldots,v_{r}\}$ and $\overline{I^{w}}=\cap_{i=1}^{r}\mathfrak{p}_{i}^{(\lceil wa_{i}\rceil)}$ for each $w\in\mathbb{Q}_{\geqslant 0}$
For a particular example, consider $R$ a polynomial ring over a field and $H_{1},\ldots,H_{t}$ subspaces of $R_{1}$. For each $i$ let $I_{i}$ be the ideal generated by $H_{i}$ and set $I=I_{1}\cdots I_{t}$. The powers of $I$ are all integrally closed and irredundant primary decompositions of them having the same format above can be found in [16, Theorem 3.2]. Therefore, these primary decompositions can be used to explicitly describe the Rees valuations and rational powers of $I$.
The following proposition shows that for $w\gg 0$, every $\overline{I^{w}}$ can be expressed in terms of finitely many rational powers of $I$.
[…]
Adopt Section missing 2 and assume $R$ is analytically unramified. Let ${\operatorname{\mathscr{RV}}}(I)=\{v_{1},\dots,v_{r}\}$ and set $e:={\operatorname{lcm}}(v_{1}(I),\dots,v_{r}(I))$ (or any multiple of it). Let $u$ be a variable of degree 1, and consider the ${\mathbb{Z}}_{\geqslant 0}$-graded $R$-algebras
$A:=R[Iu,Iu^{2},\ldots,Iu^{e}]\subseteq B:=R[u].$
If $C$ is the integral closure of $A$ in $B$, then
  1. (1)
    $C=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{\frac{n}{e}}}u^{n}$.
  2. (2)
    $C$ is a finitely generated $A$-module and furthermore a finitely generated $R$-algebra.
  3. (3)
    There exists $r,m\in{\mathbb{Z}}_{>0}$ such that $\overline{I^{mt}}=\left(\overline{I^{m}}\right)^{t}$ for every $t\in{\mathbb{Z}}_{\geqslant 0}$ and for every $w\geqslant\frac{r}{e}$ we have $\overline{I^{w+m}}=\overline{I^{w}}\,\overline{I^{m}}.$
Proof.
We begin with (1). By [38, Theorem 2.3.2], the algebra $C$ is ${\mathbb{Z}}_{\geqslant 0}$-graded and its $e$th-Veronese is the integral closure of $A_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{n}u^{en}$ in $B_{e}:=R[u^{e}]$, which is $C_{e}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{n}}u^{en}$ by [38, Proposition 5.2.1]. Write $C=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}J_{n}u^{n}$ and fix $n$. If $x\in J_{n}$, then $x^{e}\in J_{en}=\overline{I^{n}}$, and so $x\in\overline{I^{\frac{n}{e}}}$. On the other hand, if $x\in\overline{I^{\frac{n}{e}}}$, then $x^{e}\in\overline{I^{n}}$, and so $(xu^{n})^{e}\in C$. Thus, $xu^{n}$ is integral over $A$ and then $xu^{n}\in C$. It follows that $x\in J_{n}$.
We continue with (2). Note that $A=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I^{\lceil\frac{n}{e}\rceil}u^{n}$ and is Noetherian. By Rees’ theorem [51] (see also [38, Theorem 9.1.2]), there exists $k\in{\mathbb{Z}}_{>0}$ such that, for every integer $n\geqslant ek$, we have
$\overline{I^{\frac{n}{e}}}\subseteq\overline{I^{\lfloor\frac{n}{e}\rfloor}}\subseteq I^{\lfloor\frac{n}{e}\rfloor-k+1}\subseteq I^{\lceil\frac{n}{e}\rceil-k}.$
If $C_{\geqslant ek}:=\oplus_{n\geqslant ek}\overline{I^{\frac{n}{e}}}u^{n}$, then the latter inclusions show that $C_{\geqslant ek}\subset Au^{ek}$. Therefore, $C_{\geqslant ek}$, and so $C$, is a finitely generated $A$-module. In particular, $C$ is a finitely generated $R$-algebra.
Finally we prove (3). By (2) and [28, Lemma 13.10], there exists $m\in{\mathbb{Z}}_{>0}$ such that $\overline{I^{mt}}=\left(\overline{I^{m}}\right)^{t}$ for every $t\in{\mathbb{Z}}_{\geqslant 0}$. Set $C_{m}:=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\overline{I^{mt}}u^{emt}$. From (2), it follows that $C$ is a finitely generated $C_{m}$-algebra, and since $C_{m}\subset C$ is an integral extension, by [38, Lemma 2.1.9], $C$ is a finitely generated $C_{m}$-module. The statement now follows by considering $r\in{\mathbb{Z}}_{>0}$ to be the largest degree of an element in a finite set of generators of $C$ as a $C_{m}$-module. ∎
We continue with the following statement that allows for base change in rational powers (cf. [38, Corollary 19.5.2]). A ring homomorphism $f:R\to S$ is normal if it is flat and if for every $\mathfrak{p}\in{\operatorname{Spec\,}}(R)$ and field extension $\mathbb{K}$ of $R_{\mathfrak{p}}/\mathfrak{p}R_{\mathfrak{p}}$ the ring $\mathbb{K}\otimes_{R}S$ is normal.
[…]
Adopt Section missing 2. Let $f:R\to S$ be a normal homomorphism with $S$ Noetherian. For every $w\in{\mathbb{Q}}_{\geqslant 0}$ we have $\overline{I^{w}}S=\overline{(IS)^{w}}.$
Proof.
Let $e_{1}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(I)\}\right)$, $e_{2}:={\operatorname{lcm}}\left(\{v(I)\mid v\in{\operatorname{\mathscr{RV}}}(IS)\}\right)$, and $e:={\operatorname{lcm}}(e_{1},e_{2})$. Let $u,A,B,C$ be as in Section missing 2 defined using the latter $e$. By [38, Theorem 19.5.1], $C\otimes_{R}S$ is the integral closure of $A\otimes_{R}S$ in $B\otimes_{R}S$. Under the isomorphism $B\otimes_{R}S\cong S[u]$, $A\otimes_{R}S$ corresponds to $S[(IS)u,(IS)u^{2},\ldots,(IS)u^{e}]$ and $C\otimes_{R}S$ to $\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}(\overline{I^{\frac{n}{e}}}S)u^{n}$. The result follows by applying Section missing 2(1) to $IS$. ∎
Rational powers of ideal sheaves
In this subsection we introduce the notion of rational powers of ideal sheaves in normal varieties over algebraically closed fields, inspired by [45, Définition 4.2.2] and [44, Definition 9.6.2].
[…]
Let $X$ be a normal variety (integral, separated, and of finite type) over an algebraically closed field $\mathbb{k}$ and let $\mathcal{I}$ be a nonzero $\mathcal{O}_{X}$-ideal.
[…]
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor. Note that $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. The integral closure of $\mathcal{I}$, denoted by $\overline{\mathcal{I}}$, is the $\mathcal{O}_{X}$-ideal ¯$\overline{\mathcal{I}}:=v_{*}\left(\mathcal{O}_{X^{+}}(-E)\right)$ (cf. [44, Definition 9.6.2]). Moreover, for $w\in\mathbb{Q}_{\geqslant 0}$, we call the $\mathcal{O}_{X}$-ideal
$\overline{\mathcal{I}^{w}}:=v_{*}\left(\mathcal{O}_{X^{+}}(-\lceil wE\rceil)\right)$
the $w$-th rational power of $\mathcal{I}$.
[…]
If $X$ in Section missing 2 is affine, so that $\mathcal{I}=\tilde{I}$ for some $\mathbb{k}[X]$-ideal $I$, then from Section missing 2 it follows that $\overline{I^{w}}=\Gamma(X,\overline{\mathcal{I}^{w}})$ are the rational powers of $I$ as in Section missing 2.
[…]
Write $E=\sum_{i=1}^{r}e_{i}E_{i}$, where $E_{1},\ldots,E_{r}$ are distinct prime divisors of $X^{+}$, and set $e:={\operatorname{lcm}}(e_{1},\ldots,e_{r})$ (or any multiple of it). Fix $w\in\mathbb{Q}_{>0}$ and let $n:=\lceil we\rceil$. Then, $\overline{\mathcal{I}^{w}}=\overline{\mathcal{I}^{\frac{n}{e}}}$. Indeed, since the statement is local we can assume $X$ is affine and that all the $E_{i}$ are supported in $X$. Fix $x\in\Gamma\left(X,\overline{\mathcal{I}^{w}}\right)$ so that ${\operatorname{ord}}_{E_{i}}(v^{*}(x))\geqslant we_{i}$ for every $i$. Therefore, $\frac{e}{e_{i}}{\operatorname{ord}}_{E_{i}}(v^{*}(x))\geqslant n$ and it follows that ${\operatorname{ord}}_{E_{i}}(v^{*}(x))\geqslant\frac{n}{e}e_{i}$. We conclude $\overline{\mathcal{I}^{w}}\subseteq\overline{\mathcal{I}^{\frac{n}{e}}}$. Since the reverse inclusion clearly holds, we obtain $\overline{\mathcal{I}^{w}}=\overline{\mathcal{I}^{\frac{n}{e}}}$.
We have the following alternative way of computing rational powers similar to [44, Remark 9.6.4] and [45, Théorème 2.1]. We use the definition of projective morphism from [25], which is more general than the one of [34] (see [55, tag 01W7, Lemma 29.43.3]), and includes blowups of arbitrary varieties along ideal sheaves [55, tag 01OF, Lemma 31.32.13].
[…]
Adopt Section missing 2. If $v^{\prime}:X^{\prime}\longrightarrow X$ is a projective birrational morphism such that $X^{\prime}$ is a normal variety and $\mathcal{I}\cdot\mathcal{O}_{X^{\prime}}=\mathcal{O}_{X^{\prime}}(-F)$, for some effective Cartier divisor $F$ on $X^{\prime}$, then
$\overline{\mathcal{I}^{w}}=v^{\prime}_{*}\left(\mathcal{O}_{X^{\prime}}(-\lceil wF\rceil)\right).$
Proof.
By the universal properties of blowups and normalization [34, Proposition II.7.14, Exercise II.3.8], the map $v^{\prime}$ factors through the normalized blowup $v:X^{+}\longrightarrow X$. Let $g:X^{\prime}\longrightarrow X^{+}$ be the factoring map. Let $E$ be such that $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. We have the following isomorphism induced by $g^{*}\left(\mathcal{O}_{X^{+}}\right)=g^{-1}\mathcal{O}_{X^{+}}\otimes_{g^{-1},\mathcal{O}_{X^{+}}}\mathcal{O}_{X^{\prime}}\xrightarrow{\,\,\,\,\sim\,\,\,\,}\mathcal{O}_{X^{\prime}}$
(1) $g^{*}(\mathcal{O}_{X^{+}}(-E))=g^{-1}(\mathcal{O}_{X^{+}}(-E))\otimes_{g^{-1}\mathcal{O}_{X^{+}}}\mathcal{O}_{X^{\prime}}\xrightarrow{\,\,\,\,\sim\,\,\,\,}\mathcal{O}_{X^{+}}(-E)\cdot\mathcal{O}_{X^{\prime}}=\mathcal{I}\cdot\mathcal{O}_{X^{\prime}}=\mathcal{O}_{X^{\prime}}(-F).$
Indeed, a surjection of invertible sheaves is an isomorphism. Since $g$ is also projective [55, tag 01W7, Lemma 29.43.15] we have $g_{*}\left(\mathcal{O}_{X^{\prime}}\right)=\mathcal{O}_{X^{+}}$ by [34, proof of Corollary III.11.4] and [55, tag 01W7, Lemma 29.43.4]. Therefore, for every $n\in{\mathbb{Z}}_{>0}$ we have the following commutative diagram of natural maps.
[…]
It follows that $\mathcal{O}_{X^{+}}(-nE)=g_{*}\left(\mathcal{O}_{X^{\prime}}(-nF)\right)$, and this implies $v^{\prime}_{*}\left(\mathcal{O}_{X^{\prime}}(-nF)\right)=v_{*}\left(\mathcal{O}_{X^{+}}(-nE)\right)=\overline{\mathcal{I}^{n}}$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$. Set $F=\sum_{i=1}^{r}e_{i}F_{i}$, where $F_{1},\ldots,F_{r}$ are distinct prime divisors of $X^{\prime}$. Working on affine subsets of $X$, we have shown that the integral closures of the powers of $\mathcal{I}$ are defined via the valuations ${\operatorname{ord}}_{F_{i}}$. Therefore, Section missing 2 implies that the same is true for rational powers of $\mathcal{I}$, completing the proof. ∎
The previous lemma yields the following version of the birrational transformation rule for rational powers of ideals (cf. [44, Theorem 9.2.33], [45, Lemme 4.15]).
Theorem 2.2 (Birrational transformation rule for rational powers).
Adopt Section missing 2. Let $f:Y\longrightarrow X$ be a projective birational morphism with $Y$ a normal variety and set $\mathcal{I}_{Y}:=\mathcal{I}\cdot\mathcal{O}_{Y}$. For every $w\in\mathbb{Q}_{\geqslant 0}$ we have
$\overline{\mathcal{I}^{w}}=f_{*}(\overline{\mathcal{I}^{w}_{Y}}).$
Proof.
Let $\mu:Y^{+}\longrightarrow Y$ be the normalization of the blowup of $Y$ along $\mathcal{I}_{Y}$ and $D$ its exceptional divisor so that $(f\mu)^{-1}\mathcal{I}\cdot\mathcal{O}_{Y^{+}}=\mu^{-1}\mathcal{I}_{Y}\cdot\mathcal{O}_{Y^{+}}=\mathcal{O}_{Y^{+}}(-D)$. Then, by Section missing 2 we have
$\overline{\mathcal{I}^{w}}=f_{*}\mu_{*}\left(\mathcal{O}_{Y^{+}}(-\lceil wD\rceil)\right)=f_{*}(\overline{\mathcal{I}^{w}_{Y}}),$
as desired. ∎
3. Rees valuations and rational powers of invariant ideals
In this section we describe the Rees valuations and rational powers of several classes of invariant ideals. This will prove Theorem A. We begin with the following setup.
[…]
Let $\mathbb{k}$ be a field and let $R$ be a finitely generated $\mathbb{k}$-algebra. Assume $R$ is an integral domain and denote by $\operatorname{Quot}(R)$ its field of fractions. We set $I\subset R$ to be an ideal and :=$\mathcal{B}:=\{\mathfrak{b}_{\ell}\mid\ell\in\Lambda\}\subseteq R$ a $\mathbb{k}$-basis of $R$.
The following definitions will be used in the main results of this section.
[…]
Adopt Section missing 3. We say that $v:R\rightarrow{\mathbb{Z}}_{\geqslant 0}$ is a $\mathcal{B}$-monomial function (on $R$) if for every $f=\sum_{c_{i}\in\mathbb{k}^{*}}c_{i}\mathfrak{b}_{\ell_{i}}\in R$ we have $v(f)=\min\{v\left(\mathfrak{b}_{\ell_{i}}\right)\}$. If in addition $v$ is a discrete valuation of $R$, then we say that it is a $\mathcal{B}$-monomial valuation.
[…]
We say that a polyhedron $\Gamma\subseteq\mathbb{R}^{d}_{\geqslant 0}$ is positive if it is of the form $\Gamma=\mathscr{Q}+\mathbb{R}_{\geqslant 0}^{d}$ with $\mathscr{Q}\subseteq\mathbb{R}_{\geqslant 0}^{d}$ a rational polytope. We say that a (supporting) hyperplane $H$ of $\Gamma$ is non-coordinate if it is defined by $\langle\mathbf{h},\mathbf{X}\rangle=c$ where $\mathbf{h}=(h_{1},\ldots,h_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$, $c\in\mathbb{Z}_{>0}$, and $\mathbf{X}=(X_{1},\ldots,X_{d})$ are the coordinate variables of $\mathbb{R}^{d}$.
The following proposition provides a description of the Rees valuations and rational powers of ideals satisfying a natural assumption.
[…]
Adopt Section missing 3. Let $v_{1},\ldots,v_{d}$ be $\mathcal{B}$-monomial functions of $R$ and set $\underline{v}=(v_{1},\ldots,v_{d})$. Let $\Gamma\subseteq\mathbb{R}^{d}_{\geqslant 0}$ be a positive polyhedron such that for every $n\in{\mathbb{Z}}_{\geqslant 0}$ the integral closure $\overline{I^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\mathfrak{b}\in\mathcal{B}\mid\underline{v}(\mathfrak{b})\in n\Gamma\}$. Let $H_{1},\ldots,H_{r}$ be the non-coordinate hyperplanes of $\Gamma$ where $H_{k}$ is defined by $\langle\mathbf{h}_{k},\mathbf{X}\rangle=c_{k}\in{\mathbb{Z}}_{>0}$ and $\mathbf{h}_{k}=(h_{k,1},\ldots,h_{k,d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$. Assume that for each $k$ the $\mathcal{B}$-monomial function $V_{k}:=\langle\mathbf{h}_{k},\underline{v}\rangle=h_{k,1}v_{1}+\ldots+h_{k,d}v_{d}$ is a valuation of $R$. Then we have
  1. (1)
    ${\operatorname{\mathscr{RV}}}(I)=\{V_{1},\ldots,V_{r}\}$.
  2. (2)
    For every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{I^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\mathfrak{b}\in\mathcal{B}\mid\underline{v}(\mathfrak{b})\in w\Gamma\}.$
Proof.
Part (1) follows from Section missing 2. Part (2) follows from (1) since $f=\sum_{c_{i}\in\mathbb{k}^{*}}c_{i}\mathfrak{b}_{\ell_{i}}\in\overline{I^{w}}$ if and only if $V_{k}(\mathfrak{b}_{\ell_{i}})\geqslant wc_{k}$ for every $i,k$ if and only if $\underline{v}(\mathfrak{b}_{\ell_{i}})\in w\Gamma$ for every $i$. ∎
The previous proposition leads to the following notion.
[…]
If $\mathcal{B}$, $\underline{v}=(v_{1},\ldots,v_{d})$, and $\Gamma$ as in Section missing 3 exist, then we say that $(\mathcal{B},\underline{v},\Gamma)$ is a Rees package for $I$.
The following lemma offers a sufficient condition for the functions $V_{k}$ in Section missing 3 to be valuations.
[…]
Adopt Section missing 3. Let $v_{1},\ldots,v_{d}$ be $\mathcal{B}$-monomial valuations and set $\underline{v}:=(v_{1},\ldots,v_{d})$. Assume that for every $f=\sum_{c_{i}\in\mathbb{k}^{*}}c_{i}\mathfrak{b}_{\ell_{i}}$ and $g=\sum_{e_{i}\in\mathbb{k}^{*}}e_{i}\mathfrak{b}_{\mu_{i}}$ in $R$ with $\underline{v}(\mathfrak{b}_{\ell_{i}})=\underline{v}(\mathfrak{b}_{\ell_{j}})$ and $\underline{v}(\mathfrak{b}_{\mu_{i}})=\underline{v}(\mathfrak{b}_{\mu_{j}})$ for every $i,j$, there exist $\mathfrak{b}\in\mathcal{B}$ in the expansion of $fg$ as a linear combination of $\mathcal{B}$ such that $\underline{v}(\mathfrak{b})=\underline{v}(f)+\underline{v}(g)$. Then for every nonzero $\mathbf{h}=(h_{1},\ldots,h_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$ the $\mathcal{B}$-monomial function $V:=\langle\mathbf{h},\underline{v}\rangle=h_{1}v_{1}+\ldots+h_{d}v_{d}$ is a valuation of $R$.
Proof.
Fix $f=\sum_{c_{i}\in\mathbb{k}^{*}}c_{i}\mathfrak{b}_{\ell_{i}}$ and $g=\sum_{e_{i}\in\mathbb{k}^{*}}e_{i}\mathfrak{b}_{\mu_{i}}$ in $R$. Since $V$ is $\mathcal{B}$-monomial, we clearly have $V(f+g)\geqslant\min\{V(f),V(g)\}$. Notice that $V(fg)=V(f)+V(g)$ is equivalent to $V(\mathfrak{b})=V(f)+V(g)$ for some $\mathfrak{b}\in\mathcal{B}$ in the expansion of $fg$. To show the existence of such $\mathfrak{b}$ we can further assume that for every $i$ one has $V(\mathfrak{b}_{\ell_{i}})=V(f)$ and $V(\mathfrak{b}_{\mu_{i}})=V(g)$. Let $\mathscr{P}_{f}:={\operatorname{conv}}\left(\{\underline{v}(\mathfrak{b}_{\ell_{i}})\}\right)$ and $\mathscr{P}_{g}:={\operatorname{conv}}\left(\{\underline{v}(\mathfrak{b}_{\mu_{i}})\}\right)$ and notice that these polytopes belong to the parallel hyperplanes $\langle\mathbf{h},\mathbf{X}\rangle=V(f)$ and $\langle\mathbf{h},\mathbf{X}\rangle=V(g)$, respectively. Consider a vertex $\mathbf{a}$ of the Minkowski sum $\mathscr{P}_{f}+\mathscr{P}_{g}$. Therefore, there exist unique $\mathbf{a}_{f}\in\mathscr{P}_{f}\cap{\mathbb{Z}}_{\geqslant 0}^{d}$ and $\mathbf{a}_{g}\in\mathscr{P}_{g}\cap{\mathbb{Z}}_{\geqslant 0}^{d}$ such that $\mathbf{a}_{f}+\mathbf{a}_{g}=\mathbf{a}$, and then we can assume $\underline{v}(\mathfrak{b}_{\ell_{i}})=\mathbf{a}_{f}$ and $\underline{v}(\mathfrak{b}_{\mu_{j}})=\mathbf{a}_{g}$ for every $i,j$. The statement now follows by the assumption. ∎
3.1. Torus-invariant (monomial) ideals in semigroup rings
In this subsection we show that Torus-invariant ideals, i.e. those generated by monomials, in semigroup rings have a Rees package.
[…]
Let $\mathscr{S}\subset\mathbb{Z}^{s}$ be an affine semigroup of rank $s$. Let $\mathbb{k}$ be a field and let $\mathbb{k}[\mathscr{S}]$ be the semigroup ring of $\mathscr{S}$, i.e., the subalgebra of the Laurent polynomial ring $\mathbb{k}[x_{1}^{\pm 1},\dots,x_{s}^{\pm 1}]$ spanned as a vector space by the set of monomials $\mathbf{x}^{\mathscr{S}}:=\{\mathbf{x}^{\mathbf{n}}=x_{1}^{n_{1}}\cdots x_{s}^{n_{s}}\mid\mathbf{n}=(n_{1},\ldots,n_{s})\in\mathscr{S}\}$, and multiplication induced by the operation of $\mathscr{S}$. Let $\mathcal{C}(\mathscr{S}):=\mathbb{R}_{\geqslant 0}\mathscr{S}$ be the cone generated by $\mathscr{S}$ and assume it is strongly convex, i.e., $\mathcal{C}(\mathscr{S})$ does not contain any subspace of positive dimension. Let $\mathcal{F}_{1},\ldots,\mathcal{F}_{d}$ be the facets of $\mathcal{C}(\mathscr{S})$ and for each $k$ let $\mathcal{F}_{k}$ be defined by $\langle\mathbf{f}_{k},\mathbf{X}\rangle=0$ with $\mathbf{f}_{k}\in\mathbb{Z}^{s}$. Let $v_{k}$ be the $\mathbf{x}^{\mathscr{S}}$-monomial valuation determined by $\mathbf{x}^{\mathbf{n}}\mapsto\langle\mathbf{f}_{k},\mathbf{n}\rangle$.
Let $I\subseteq\mathbb{k}[\mathscr{S}]$ be a monomial ideal, i.e., it is generated by monomials. Set $\log(I):=\{\mathbf{n}\in\mathscr{S}\mid\mathbf{x}^{\mathbf{n}}\in I\}$ and $\Gamma(I):={\operatorname{conv}}\left(\log(I)\right)\subseteq\mathbb{R}^{s}$.
We refer to §missing 3.1 for a visual representation of the previous notations. The following proposition is an important ingredient to describe the Rees package of monomial ideals.
[…]
Adopt §missing 3.1. The integral closure $\overline{I}$ is a monomial ideal generated by the monomials $\{\mathbf{x}^{\mathbf{n}}\mid\mathbf{n}\in\Gamma(I)\cap\mathscr{S}\}$.
Proof.
Let $\tilde{\mathscr{S}}:=\mathbb{Z}\mathscr{S}\cap\mathbb{R}_{\geqslant 0}\mathscr{S}$ be the normalization of $\mathscr{S}$. By [9, Theorem 4.46] the statement holds for $I\,\mathbb{k}[\tilde{\mathscr{S}}]$. The statement now follows for $I$ by [38, Proposition 1.6.1] since $\mathbb{k}[\mathscr{S}]\subseteq\mathbb{k}[\tilde{\mathscr{S}}]$ is an integral extension [9, Lemma 4.41] and $\Gamma(I\,\mathbb{k}[\tilde{\mathscr{S}}])=\Gamma(I)$. ∎
We are now ready to describe a Rees package for monomial ideals.
Theorem 3.1.
Adopt §missing 3.1 . If $\underline{v}:=(v_{1},\ldots,v_{d})$ and $\Sigma:={\operatorname{conv}}\left(\underline{v}\left(\log(I)\right)\right)\subseteq\mathbb{R}_{\geqslant 0}^{d}$, then $(\mathbf{x}^{\mathscr{S}},\underline{v},\Sigma)$ is a Rees package for $I$. In particular, the Rees valuations of $I$ are determined by the non-coordinates hyperplanes of $\Sigma$ and for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{I^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\mathbf{x}^{\mathbf{n}}\in\mathbf{x}^{\mathscr{S}}\mid\underline{v}\left(\mathbf{x}^{\mathbf{n}}\right)\in w\Sigma\}=\{\mathbf{x}^{\mathbf{n}}\mid\mathbf{n}\in w\Gamma(I)\cap\mathscr{S}\}$.
Proof.
Clearly $\Gamma(I^{n})=n\Gamma(I)$ for every $n\in{\mathbb{Z}}_{\geqslant 0}$, and so $n\Sigma={\operatorname{conv}}\left(\underline{v}\left(\log(I^{n})\right)\right)$. Also, for every $\mathbf{m}\in{\mathbb{Z}}_{\geqslant 0}^{d}$ the set $\underline{v}^{-1}(\mathbf{m})\cap\mathbf{x}^{\mathscr{S}}$ has at most one element, therefore the assumption in Section missing 3 is trivially satisfied by $\mathbf{x}^{\mathscr{S}}$ and $\underline{v}$. The statements then follow from §missing 3.1, Section missing 3, and Section missing 3. ∎
[…]
We note that Theorem Theorem 3.1 applies to any ideal whose integral closure is monomial. Such ideals are called Newton non-degenerate in the literature; see [53].
In the next example we describe the notation in this section for one specific ideal.
[…]
We follow the notation from Theorem 3.1. Let $\mathscr{S}\subset\mathbb{Z}^{2}$ be the (normal) semigroup generated by $\{(2,1),(1,3)\}$, see Figure 1 below. Thus $\mathbb{k}[\mathscr{S}]=\mathbb{k}[x_{1}^{2}x_{2},x_{1}x_{2}^{3}]$. The valuations $v_{1},v_{2}$ corresponding to the rays $\mathcal{F}_{1},\mathcal{F}_{2}$ are determined by $\mathbf{x}^{\mathbf{n}}\mapsto-n_{1}+2n_{2}$ and $\mathbf{x}^{\mathbf{n}}\mapsto 3n_{1}-n_{2}$, respectively. The ideal $I\subset\mathbb{k}[\mathscr{S}]$ is generated by $x_{1}^{4}x_{2}^{2}$ and $x_{1}^{3}x_{2}^{4}$. Therefore $\Gamma(I)$ has as facets the rays $(4,2)+\mathbb{R}_{\geqslant 0}(2,1)$, $(3,4)+\mathbb{R}_{\geqslant 0}(1,3)$, and the segment ${\operatorname{conv}}\left((3,4),(4,2)\right)$. It follows that $\Sigma$ has supporting hyperplanes $H:v_{1}=0$, $H_{1}:v_{1}+v_{2}=10$, and $H_{2}:v_{2}=5$. The latter two, $H_{1},H_{2}$, are the non-coordinate ones and so these are the ones that determine the Rees valuations of $I$.
Figure 1. The semigroup $\mathscr{S}={\mathbb{Z}}_{\geqslant 0}(2,1)+{\mathbb{Z}}_{\geqslant 0}(1,3)$ and ideal $I=\left(x_{1}^{4}x_{2}^{2},\,x_{1}^{3}x_{2}^{4}\right)\subset\mathbb{k}[\mathscr{S}]$.
3.2. Sums of products of determinantal ideals and $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals
In this subsection we show that ideals obtained by taking sums of products of determinantal ideals in a generic matrix $X$ have a Rees package. As a consequence we obtain that $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of the polynomial ring $\mathbb{C}[X]$ have a Rees package. Throughout we follow the treatment in [17, 8, 35, 6].
[…]
Let $X$ be an $m\times n$ generic matrix with $m\leqslant n$. Let $\mathbb{k}$ be a field and $R=\mathbb{k}[X]$ the corresponding polynomial ring. For any $t\leqslant m$ let $I_{t}\subset R$ denote the ideal generated by the $t$-minors of $X$. The $t$-minor corresponding to rows $i_{1}<\cdots<i_{t}\leqslant m$ and columns $j_{1}<\cdots<j_{t}\leqslant n$ will be denoted as $[i_{1},\dots,i_{t}|j_{1},\dots,j_{t}]$. For an positive integer $p$ we denote by $[p]$ the set $[p]:=\{1,\ldots,p\}$. A diagram $\sigma$ is a vector $\sigma:=(s_{1},\ldots,s_{p})\in{\mathbb{Z}}_{\geqslant 0}^{p}$ with $s_{1}\geqslant\cdots\geqslant s_{p}$. If $\tau=(t_{1},\ldots,t_{q})$ and $\sigma=(s_{1},\ldots,s_{p})$ are two diagrams, we write $\tau\subseteq\sigma$ if $q\leqslant p$ and $t_{i}\leqslant s_{i}$ for every $i\in[q]$. Let $\mathcal{H}_{m}$ be the set of all diagrams such that $s_{i}\leqslant m$ for every $i$. For $t\in[m]$ we denote by $\gamma_{t}$ the function $\gamma_{t}:\mathcal{H}_{m}\to{\mathbb{Z}}_{\geqslant 0}$ given by $\gamma_{t}(s_{1},\ldots,s_{p}):=\sum_{i=1}^{p}\max\{0,s_{i}-t+1\}$.
We define a partial order on the minors of $X$ in the following way:
$[i_{1},\dots,i_{t}|j_{1},\dots,j_{t}]\leqslant[e_{1},\dots,e_{h}|f_{1},\dots,f_{h}]\iff t\geqslant h\text{ and }i_{q}\leqslant e_{q},\,j_{q}\leqslant~{}f_{q}~{}\text{ for every }q\in[h].$
A standard monomial is a product of minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{m}$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [17, Theorem 2.1] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each diagram $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{m}$ we denote by $I_{\sigma}$ the product of determinantal ideals $I_{\sigma}:=I_{s_{1}}\cdots I_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{m}$ we set
$I_{\Lambda}:=\sum_{\sigma\in\Lambda}I_{\sigma}.$
For each $t\in[m]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [17, Theorem 7.1].
We now show that the ideals $I_{\Lambda}$ have a Rees package. In the following, by $I^{(n)}$ we denote the $n$th-symbolic power of the ideal $I$.
Theorem 3.2.
Adopt §missing 3.2 and the notation above. Let $\Lambda\subseteq\mathcal{H}_{m}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{m})$ and $\Gamma:={\operatorname{conv}}\big{(}\underline{\gamma}(\Lambda)\big{)}+\mathbb{R}_{\geqslant 0}^{m}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $I_{\Lambda}$. In particular, the Rees valuations of $I_{\Lambda}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{I_{\Lambda}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
$\overline{I_{\Lambda}^{w}}=\sum_{(a_{1},\dots,a_{m})\in w\Gamma\cap{\mathbb{Z}}_{\geqslant 0}^{m}}\left(\bigcap_{i=1}^{m}I_{i}^{(a_{i})}\right).$
Proof.
By [35, Theorems 2.2 & 2.3] the integral closure $\overline{I_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By [8, Proposition 7.4.2, Equations (7.2) & (7.3)] the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$; indeed, notice that by [8, Remark 3.2.9] we may assume $\mathbb{k}=\overline{\mathbb{k}}$, and under this assumption Segre products of domains are domains. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $I_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows from [35, Theorems 2.2 & 2.3]. ∎
In the next example we illustrate the previous theorem.
[…]
Let $X$ be a $2\times 3$ generic matrix and $\Lambda=\{(2),\,(1,1,1)\}$, i.e., $I_{\Lambda}=I_{2}+I_{1}^{3}$. Since $\underline{\gamma}(2)=(2,1)$ and $\underline{\gamma}(1,1,1)=(3,0)$ we have $\Gamma={\operatorname{conv}}\left((2,1),(3,0)\right)+\mathbb{R}^{2}_{\geqslant 0}$. The non-coordinate hyperplanes of $\Gamma$ are $X_{1}=2$ and $X_{1}+X_{2}=3$, see Figure 2. Therefore the Rees valuations of $I_{\Lambda}$ are $\gamma_{1}$ and $\gamma_{1}+\gamma_{2}$. Finally, for every $w$ we have
$\overline{I_{\Lambda}^{w}}=\sum_{\stackrel{{\scriptstyle a_{1},a_{2}\in{\mathbb{Z}}_{\geqslant 0}}}{{a_{1}\geqslant 2w,a_{1}+a_{2}\geqslant 3w}}}I_{1}^{(a_{1})}\cap I_{2}^{(a_{2})}.$
Figure 2. The polyhedron $\Gamma={\operatorname{conv}}\big{(}\underline{\gamma}(\Lambda)\big{)}+\mathbb{R}_{\geqslant 0}^{2}={\operatorname{conv}}\left((2,1),(3,0)\right)+\mathbb{R}^{2}_{\geqslant 0}$.
[…]
As defined in [30], the asymptotic resurgence of a homogeneous ideal $I$ in a polynomial ring is $\rho_{a}(I):=\sup\{s/r\mid I^{(st)}\not\subset I^{rt}\text{ for all }t\gg 0\}$. In [26, Theorem 4.10] the authors showed that $\rho_{a}(I)=\max_{v\in{\operatorname{\mathscr{RV}}}(I)}\left\{v(I)/\widehat{v}(I)\right\}$, where $\widehat{v}(I)=\lim_{r\to\infty}v(I^{(r)})/r$. By Theorem 3.2, if $I_{t}$ is the determinantal ideal defined in §missing 3.2, then ${\operatorname{\mathscr{RV}}}(I_{t})=\{\gamma_{1},\ldots,\gamma_{t}\}$. Using [17, Theorem 7.1] and [22, Remark 7.23] one shows that $\rho_{a}(I_{t})=\dfrac{\gamma_{1}(I_{t})}{\displaystyle\widehat{\gamma_{1}}(I_{t})}=\dfrac{t(m-t+1)}{m}$.
For the rest of this subsection assume $\mathbb{k}=\mathbb{C}$ and set $G:=\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$. Consider the action of $G$ on $R=\mathbb{C}[X]$ induced by $X_{i,j}\mapsto(AXB^{-1})_{i,j}$ for $(A,B)\in G$. In [17, Theorem 5.1] the authors described certain irreducible representations $L_{\sigma}$ and $\prescript{}{\sigma}{L}$ of $\operatorname{GL}_{m}(\mathbb{C})$ and $\operatorname{GL}_{n}(\mathbb{C})$, respectively, indexed by $\sigma\in\mathcal{H}_{m}$, and they showed that an ideal of $R$ is invariant under the action of $G$ if and only if it is of the form
$\mathcal{I}(\Lambda):=\bigoplus_{\stackrel{{\scriptstyle\tau\subseteq\sigma}}{{\text{for some }\tau\in\Lambda}}}L_{\sigma}\otimes_{\mathbb{C}}\prescript{}{\sigma}{L}$
for some subset $\Lambda\subseteq\mathcal{H}_{m}$. Moreover, in [17, Theorem 8.2] it is shown that $\overline{\mathcal{I}(\Lambda)}=\overline{I_{\Lambda}}$. Thus we obtain the following corollary.
[…]
Under the assumptions and notations in Theorem 3.2 with $\mathbb{k}=\mathbb{C}$, for every $\Lambda\subseteq\mathcal{H}_{m}$ the $G$-invariant ideal $\mathcal{I}(\Lambda)$ has a Rees package.
3.3. Sums of products of determinantal ideals of symmetric matrices and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals
In this subsection we show that ideals obtained by taking sums of products of determinantal ideals in a symmetric generic matrix $Y$ have a Rees package. As a consequence we obtain that $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of the polynomial ring $\mathbb{C}[Y]$ have a Rees package. Throughout we follow the treatment in [2, 35, 40].
[…]
Let $Y$ be an $n\times n$ generic symmetric matrix. Let $\mathbb{k}$ be a field and $R=\mathbb{k}[Y]$ the corresponding polynomial ring. For any $t\leqslant n$ let $J_{t}\subset R$ denote the ideal generated by the $t$-minors of $Y$. We follow the notation introduced in §missing 3.2 for minors and diagrams.
A minor $[i_{1},\dots,i_{t}|j_{1},\dots,j_{t}]$ of $Y$ is doset if $i_{q}\leqslant j_{q}$ for every $q\in[t]$. We define a partial order on the doset minors in the following way:
$[i_{1},\dots,i_{t}|j_{1},\dots,j_{t}]\leqslant[e_{1},\dots,e_{h}|f_{1},\dots,f_{h}]\iff t\geqslant h\text{ and }j_{q}\leqslant e_{q},~{}\text{ for every }q\in[h].$
A standard monomial is a product of doset minors $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{n}$ where $\delta_{i}$ is a $s_{i}$-minor and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [40, §4.2] (see also [2]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{n}$ we denote by $J_{\sigma}$ the product of determinantal ideals $J_{\sigma}:=J_{s_{1}}\cdots J_{s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{n}$ we set
$J_{\Lambda}:=\sum_{\sigma\in\Lambda}J_{\sigma}.$
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [2, Theorem 5.1], or [35, Theorem 2.7].
[…]
Adopt §missing 3.3 and the notation above. Set $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$. For every $\mathbf{n}\in{\mathbb{Z}}_{\geqslant 0}^{n}$ let $V_{\mathbf{n}}$ (resp. $V_{>\mathbf{n}}$) be the $\mathbb{k}$-subspace of $R$ spanned by the standard monomials $\underline{\delta}$ such that $\underline{\gamma}(\underline{\delta})=\mathbf{n}$ (resp. $\underline{\gamma}(\underline{\delta})>\mathbf{n}$ componentwise). Then the associated graded ring $\mathcal{G}=\oplus_{\mathbf{n}\in{\mathbb{Z}}_{\geqslant 0}^{n}}V_{\mathbf{n}}/V_{>\mathbf{n}}$ is a domain.
Proof.
Let $X$ be a generic $n\times n$ matrix and $S$ be the subalgebra of $\mathbb{k}[X]$ generated by the products of minors $[1,\ldots,t_{1}|i_{1},\ldots,i_{t_{1}}][1,\ldots,t_{2}|i_{1},\ldots,i_{t_{2}}]$ for some $1\leqslant i_{1},i_{2}\leqslant n$. One proceeds as in [11, proof of Thorem 1.6] via [18, Lemma 5.2 and proof of Lemma 5.3] to show $\mathcal{G}\cong S$. The conclusion follows. ∎
We now show that the ideals $J_{\Lambda}$ have a Rees package.
Theorem 3.3.
Adopt §missing 3.3 and the notation above. Let $\Lambda\subseteq\mathcal{H}_{n}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$ and $\Gamma:={\operatorname{conv}}\big{(}\underline{\gamma}(\Lambda)\big{)}+\mathbb{R}_{\geqslant 0}^{n}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $J_{\Lambda}$. In particular, the Rees valuations of $J_{\Lambda}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{J_{\Lambda}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
$\overline{J_{\Lambda}^{w}}=\sum_{(a_{1},\dots,a_{n})\in w\Gamma\cap{\mathbb{Z}}_{\geqslant 0}^{n}}\left(\bigcap_{i=1}^{n}J_{i}^{(a_{i})}\right).$
Proof.
By [35, Theorems 2.6 & 2.7] the integral closure $\overline{J_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By §missing 3.3 the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $J_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows again from [35, Theorems 2.6 & 2.7]. ∎
For the rest of this subsection assume $\mathbb{k}=\mathbb{C}$ and set $G:=\operatorname{GL}_{n}(\mathbb{C})$. Consider the action of $G$ on $R=\mathbb{C}[Y]$ induced by $Y_{i,j}\mapsto(AYA^{t})_{i,j}$ for $A\in G$. In [2] the author described certain irreducible representations $L_{\sigma}$ of $G$ indexed by $\sigma\in\mathcal{H}_{n}$, and they showed that an ideal of $R$ is invariant under the action of $G$ if and only if it is of the form
$\mathcal{J}(\Lambda):=\bigoplus_{\stackrel{{\scriptstyle\tau\subseteq\sigma}}{{\text{for some }\tau\in\Lambda}}}L_{\sigma}$
for some subset $\Lambda\subseteq\mathcal{H}_{n}$. Moreover, in [2, Theorem 6.1] (see also [35, Theorem 2.11]) it is shown that $\overline{\mathcal{J}(\Lambda)}=\overline{J_{\Lambda}}$. Thus we obtain the following corollary.
[…]
Under the assumptions and notations in Theorem 3.3 with $\mathbb{k}=\mathbb{C}$, for every $\Lambda\subseteq\mathcal{H}_{n}$ the $G$-invariant ideal $\mathcal{J}(\Lambda)$ has a Rees package.
3.4. Sums of products of ideals of Pfaffians and $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals
In this subsection we show that ideals obtained by taking sums of products of ideals of Pfaffians in a skew-symmetric generic matrix $Z$ have a Rees package. As a consequence we obtain that $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of the polynomial ring $\mathbb{C}[Z]$ have a Rees package. Throughout we follow the treatment in [1, 35, 40].
[…]
Let $Z$ be an $n\times n$ skew-symmetric matrix of variables ($Z=-Z^{t}$). We follow the notation introduced in §missing 3.2 for minors and diagrams. Let $\mathbb{k}$ be a field and $R=\mathbb{k}[Z]$ the corresponding polynomial ring. For any $t\leqslant\lfloor n/2\rfloor$ a minor of $Z$ of the form $[i_{1},\dots,i_{2t}|i_{1},\dots,i_{2t}]$ is a square of a polynomial which is called a $2t$-Pfaffian. We will denote this Pfaffian as $[i_{1},\dots,i_{2t}]$. Let $P_{2t}\subset R$ be the ideal generated by the $2t$-Pfaffians of $Z$.
We define a partial order on Pfaffians in the following way:
$[i_{1},\dots,i_{2t}]\leqslant[e_{1},\dots,e_{2h}]\iff t\geqslant h\text{ and }i_{q}\leqslant e_{q},~{}\text{ for every }q\in[2h].$
A standard monomial in this case is a product of Pfaffians $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{1}\leqslant\dots\leqslant\delta_{p}$. For every $\underline{\delta}$ we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor n/2\rfloor}$ where $\delta_{i}$ is a $2s_{i}$-Pfaffian and call $\sigma$ the shape of $\underline{\delta}$. Let $\mathcal{B}$ be the set of all standard monomials. By [40, §4.3] (see also [1]) $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor n/2\rfloor}$ we denote by $J_{\sigma}$ the product of ideals of Pfaffians $P_{\sigma}:=P_{2s_{1}}\cdots P_{2s_{p}}$, and for any subset $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$ we set
$P_{\Lambda}:=\sum_{\sigma\in\Lambda}P_{\sigma}.$
For each $t\in\big{[}\lfloor n/2\rfloor\big{]}$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [1, Theorem 5.1], or [35, Theorem 2.10].
We now show that the ideals $P_{\Lambda}$ have a Rees package.
Theorem 3.4.
Adopt §missing 3.4 and the notation above. Let $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{\lfloor n/2\rfloor})$ and $\Gamma:={\operatorname{conv}}\big{(}\underline{\gamma}(\Lambda)\big{)}+\mathbb{R}_{\geqslant 0}^{\lfloor n/2\rfloor}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $P_{\Lambda}$. In particular, the Rees valuations of $P_{\Lambda}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{P_{\Lambda}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
$\overline{P_{\Lambda}^{w}}=\sum_{\left(a_{1},\dots,a_{\lfloor n/2\rfloor}\right)\in w\Gamma\cap{\mathbb{Z}}_{\geqslant 0}^{\lfloor n/2\rfloor}}\left(\bigcap_{i=1}^{\lfloor n/2\rfloor}P_{2i}^{(a_{i})}\right).$
Proof.
By [35, Theorems 2.9 & 2.10] the integral closure $\overline{P_{\Lambda}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. By [11, Theorem 1.6 and its proof] (see [18, Lemma 6.2 & proof of Theorem 6.5] to remove the assumption on the characteristic) the assumption in Section missing 3 is satisfied by $\mathcal{B}$ and $\underline{\gamma}$. Therefore, $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $P_{\Lambda}$ by Section missing 3 and Section missing 3. The last formula follows again from [35, Theorems 2.9 & 2.10]. ∎
For the rest of this subsection assume $\mathbb{k}=\mathbb{C}$ and set $G:=\operatorname{GL}_{n}(\mathbb{C})$. Consider the action of $G$ on $R=\mathbb{C}[Z]$ induced by $Z_{i,j}\mapsto(AZA^{t})_{i,j}$ for $A\in G$. In [1] the authors described certain irreducible representations $L_{\sigma}$ of $G$ indexed by $\sigma\in\mathcal{H}_{\lfloor n/2\rfloor}$, and they showed that an ideal of $R$ is invariant under the action of $G$ if and only if it is of the form
$\mathcal{P}(\Lambda):=\bigoplus_{\stackrel{{\scriptstyle\tau\subseteq\sigma}}{{\text{for some }\tau\in\Lambda}}}L_{\sigma}$
for some subset $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$. Moreover, in [1, Theorem 6.2] it is shown that $\overline{\mathcal{P}(\Lambda)}=\overline{P_{\Lambda}}$. Thus we obtain the following corollary.
[…]
Under the assumptions and notations in Theorem 3.4 with $\mathbb{k}=\mathbb{C}$, for every $\Lambda\subseteq\mathcal{H}_{\lfloor n/2\rfloor}$ the $G$-invariant ideal $\mathcal{P}(\Lambda)$ has a Rees package.
3.5. Products of determinantal ideals of Hankel matrices
In this subsection we show that products of determinantal ideals of Hankel matrices have a Rees package. Throughout we follow the treatment in [15].
[…]
Let $\mathbb{k}$ be a field and $R=\mathbb{k}[\underline{x}]=\mathbb{k}[x_{1},\ldots,x_{n}]$ a polynomial ring. Fix $c\in[n]$ and let $W$ be the $c\times(n-c+1)$ Hankel matrix
$W:=W_{c}=\begin{pmatrix}x_{1}&x_{2}&x_{3}&\dots&x_{n-c+1}\\ x_{2}&x_{3}&\dots&\dots&\dots\\ x_{3}&\dots&\dots&\dots&\dots\\ \vdots&\vdots&\vdots&\vdots&\vdots\\ x_{c}&\dots&\dots&\dots&x_{n}\end{pmatrix}.$
For any $t\leqslant\min\{c,n-c+1\}$ let $H_{t}\subset R$ denote the ideal generated by the $t$-minors of $W$. The ideals $H_{t}$ only depends on $t$ and $n$ [15, Corollary 2.2]. Note that the maximum value for $t$ is $\lfloor(n+1)/2\rfloor$. We follow the notation introduced in §missing 3.2 for minors and diagrams.
A minor of $W$ of the form $[1,\dots,t|a_{1},\dots,a_{t}]$ is a maximal minor and it will be denoted by $[a_{1},\dots,a_{t}]$. A monomial $x_{a_{1}}\cdots x_{a_{t}}$ of $R$ is a $<_{1}$-chain if $a_{1}<_{1}a_{2}\dots<_{1}a_{s}$, where $a<_{1}b$ means $a+1<b$. There is a bijective map $\left\{<_{1}\text{chains of }R\right\}\leftrightarrow\{\text{maximal minors}\}$ defined by $\phi(x_{a_{1}}\dots x_{a_{t}})=[a_{1},a_{2}-1,\dots,a_{t}-t+1]$ which induces an injective map
Φ:{monomials of $\Phi:\{\text{monomials of }R\}\to\{\text{product of maximal minors}\};$
see [15] for details. The elements in the image $\mathcal{B}$ of $\Phi$ are the standard monomials of $W$. By [15, Theorem 1.2] $\mathcal{B}$ forms a $\mathbb{k}$-basis for $R$. For every standard monomial $\underline{\delta}:=\delta_{1}\cdots\delta_{p}$ with $\delta_{i}$ a $s_{i}$-maximal-minor and $s_{1}\geqslant\cdots\geqslant s_{p}$, we associate a diagram $|\underline{\delta}|:=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. For each $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$ we denote by $H_{\sigma}$ the product of determinantal ideals
$H_{\sigma}:=H_{s_{1}}\cdots H_{s_{p}}.$
For each $t\in[n]$ we extend $\gamma_{t}$ to a $\mathcal{B}$-monomial valuation induced by $\gamma_{t}(\underline{\delta}):=\gamma_{t}(|\underline{\delta}|)$. Indeed these are valuations by [15, Theorem 3.8].
We now show that the ideals $H_{\sigma}$ have a Rees package.
Theorem 3.5.
Adopt §missing 3.5 and the notation above. Let $\sigma=(s_{1},\ldots,s_{p})\in\mathcal{H}_{\lfloor(n+1)/2\rfloor}$. If $\underline{\gamma}:=(\gamma_{1},\ldots,\gamma_{n})$ and $\Gamma:=\underline{\gamma}(\sigma)+\mathbb{R}_{\geqslant 0}^{n}$, then $(\mathcal{B},\underline{\gamma},\Gamma)$ is a Rees package for $H_{\sigma}$. In particular, the Rees valuations of $H_{\sigma}$ are determined by the non-coordinates hyperplanes of $\Gamma$ and ${\operatorname{\mathscr{RV}}}(H_{\sigma})\subseteq\{\gamma_{1},\ldots,\gamma_{s_{1}}\}$. Furthermore, for every $w\in{\mathbb{Q}}_{\geqslant 0}$, $\overline{H_{\sigma}^{w}}$ is spanned as a $\mathbb{k}$-vector space by $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in w\Gamma\}$ and
$\overline{H_{\sigma}^{w}}=\bigcap_{i=1}^{n}H_{i}^{(\lceil w\gamma_{i}(\sigma)\rceil)}.$
Proof.
By [15, Theorem 3.12] the integral closure $\overline{H_{\sigma}^{n}}$ is spanned as a $\mathbb{k}$-vector space by the set $\{\underline{\delta}\in\mathcal{B}\mid\underline{\gamma}(\underline{\delta})\in n\Gamma\}$. Thus the statements follow form Section missing 3 and Section missing 3. The last formula follows from [15, Theorems 3.8 & 3.12]. ∎
4. Rees package and the summation formula
This section is devoted to the proof of Theorem 4.1, which establishes that two ideals with a Rees package satisfy the summation formula for rational powers for every $w\in\mathbb{Q}_{\geqslant 0}$. From this theorem Theorem B (or Section missing 4) directly follows. We will continue using the notations and definitions in Section missing 3.
We begin with the following basic convex geometric fact. We include a sketch of its proof as we could no find a reference for it, but first we need the following notation.
[…]
Let $\Gamma_{1}\subsetneq\mathbb{R}^{n_{1}}_{\geqslant 0}$ and $\Gamma_{2}\subsetneq\mathbb{R}^{n_{2}}_{\geqslant 0}$ be two positive polyhedra. We denote by ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})\subseteq{\mathbb{Z}}_{\geqslant 0}^{n_{1}+n_{2}}$ the convex hull of $\Gamma_{1}$ and $\Gamma_{2}$ after we embed them in ${\mathbb{Z}}_{\geqslant 0}^{n_{1}+n_{2}}$ by adding zeroes in the last $n_{2}$ components and the first $n_{1}$ components, respectively. For two hyperplanes $H_{1},H_{2}$ defined by $\langle\mathbf{h}_{1},\mathbf{X}_{1}\rangle=h_{1,1}X_{1,1}+\cdots+h_{1,n_{1}}X_{1,n_{1}}=c_{1}$ and $\langle\mathbf{h}_{2},\mathbf{X}_{2}\rangle=h_{2,1}X_{2,1}+\cdots+h_{2,n_{2}}X_{2,n_{2}}=c_{2}$, respectively, we denote by $H_{1}\star H_{2}$ the hyperplane $H_{1}\star H_{2}=\langle c_{2}\mathbf{h}_{1}+c_{1}\mathbf{h}_{2},(\mathbf{X}_{1},\mathbf{X}_{2})\rangle=c_{1}c_{2}$.
[…]
Adopt Section missing 4. ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$ is a positive polyhedron and its non-coordinate supporting hyperplanes are those of the form $H_{1}\star H_{2}$ where $H_{1}$ and $H_{2}$ are non-coordinate supporting hyperplanes of $\Gamma_{1}$ and $\Gamma_{2}$, respectively.
Proof.
Write $\Gamma_{1}=\mathscr{Q}_{1}+\mathbb{R}_{\geqslant 0}^{n_{1}}$ and $\Gamma_{2}=\mathscr{Q}_{2}+\mathbb{R}_{\geqslant 0}^{n_{2}}$ where $\mathscr{Q}_{1}$ and $\mathscr{Q}_{2}$ are rational polytopes not containing the origin. By [48, Proposition 2.3] the second statement holds for ${\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})$. By [9, Proposition 1.31] and its proof we have ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})={\operatorname{conv}}(\mathscr{Q}_{1},\mathscr{Q}_{2})+\mathbb{R}_{\geqslant 0}^{n_{1}+n_{2}}$, and from this both statements follow for ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$. ∎
[…]
Consider the positive polyhedra $\Gamma_{1}=(4)+\mathbb{R}_{\geqslant 0}^{1}$ and $\Gamma_{2}={\operatorname{conv}}\left((2,1),(3,0)\right)+\mathbb{R}^{2}_{\geqslant 0}$ (cf. §missing 3.2). The only non-coordinate hyperplane $H$ of $\Gamma_{1}$ has equation $X=4$, and the ones of $\Gamma_{2}$, denoted by $H_{1}$ and $H_{2}$, have equations $X_{1}+X_{2}=3$ and $X_{1}=2$. Therefore, from Section missing 4 it follows that ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$ have two non-coordinate hyperplanes $H\star H_{1}$ and $H\star H_{2}$ with equations $124X_{1}+4X_{2}+3X=$ 12 and $42X_{1}+X=$ 4, see Figure 3.
Figure 3. The convex hull ${\operatorname{conv}}(\Gamma_{1},\Gamma_{2})$.
The following setup will be used in the rest of this section.
[…]
Let $\mathbb{k}$ be a field. Let $R$ and $S$ be finitely generated $\mathbb{k}$-algebras that are domains. Let $I\subset R$ and $J\subset S$ be nonzero ideals that have a Rees package (see Section missing 3). Let $(\mathcal{B},\underline{v}=(v_{1},\ldots,v_{d}),\Gamma)$ and $(\mathcal{C},\underline{\gamma}=(\gamma_{1},\ldots,\gamma_{e}),\Sigma)$ be Rees packages for $I$ and $J$, respectively. We denote by $\mathcal{B}\otimes\mathcal{C}$ the basis $\{\mathfrak{b}\otimes\mathfrak{c}\mid\mathfrak{b}\in\mathcal{B},\mathfrak{c}\in\mathcal{C}\}$ of $T:=R\otimes_{\mathbb{k}}S$.
The following are two technical lemmas that are needed in the main result of this section.
[…]
Adopt Section missing 4. Let $v$ and $\gamma$ be discrete valuations of $R$ and $S$, respectively, such that $v$ is $\mathcal{B}$-monomial and $\gamma$ is $\mathcal{C}$-monomial. Let $v+\gamma:T\to{\mathbb{Z}}_{\geqslant 0}$ be the $\mathcal{B}\otimes\mathcal{C}$-monomial function induced by $(v+\gamma)(\mathfrak{b}\otimes\mathfrak{c})=v(\mathfrak{b})+\gamma(\mathfrak{c})$. Then $v+\gamma$ is a valuation.
Proof.
Set $V:=v+\gamma$. It suffices to show that for every $f,g\in T$ one has $V(fg)=V(f)+V(g)$. Write
$f=\sum_{c_{i}\in\mathbb{k}^{*}}c_{i}\,\mathfrak{b}_{i}\otimes\mathfrak{c}_{i}\,\,\,\quad\text{and}\quad\,\,\,g=\sum_{d_{j}\in\mathbb{k}^{*}}d_{j}\,\mathfrak{b}_{j}\otimes\mathfrak{c}_{j}.$
Set $a:=V(f)$ and $b:=V(g)$. We may assume $V(\mathfrak{b}_{i}\otimes\mathfrak{c}_{i})=a$ and $V(\mathfrak{b}_{j}\otimes\mathfrak{c}_{j})=b$ for every $i,j$. Assume $V(fg)>a+b$, we will arrive to a contradiction. The latter means that all the elements of $\mathcal{B}\otimes\mathcal{C}$ of degree $a+b$ appearing in the expansions of $(\mathfrak{b}_{i}\otimes\mathfrak{c}_{i})(\mathfrak{b}_{j}\otimes\mathfrak{c}_{j})$ cancel in $fg$. Since these monomials would still cancel if one passes to the algebraic closure of $\mathbb{k}$, we may assume $\mathbb{k}=\overline{\mathbb{k}}$.
For each $n\in{\mathbb{Z}}_{\geqslant 0}$ consider the following $R$-ideals $I_{n}:=(\{\mathfrak{b}\in\mathcal{B}\mid v(\mathfrak{b})\geqslant n\})$ and $I_{>n}:=(\{\mathfrak{b}\in\mathcal{B}\mid v(\mathfrak{b})>n\})$, and the $S$-ideals $J_{n}:=(\{\mathfrak{c}\in\mathcal{C}\mid\gamma(\mathfrak{c})\geqslant n\})$ and $J_{>n}:=(\{\mathfrak{c}\in\mathcal{C}\mid\gamma(\mathfrak{c})>n\})$. Consider the $\mathbb{k}$-algebras $G_{1}=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}I_{n}/I_{>n}$ and $G_{2}=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}J_{n}/J_{>n}$. One easily shows that $G_{1}$ and $G_{2}$ are domains. Let $G:=G_{1}\otimes_{\mathbb{k}}G_{2}$ which is also a domain by [5, Chapter V, §17.2, Corollary]. Consider $G$ as an ${\mathbb{Z}}_{\geqslant 0}$-graded $\mathbb{k}$-algebra with $[G]_{m}=\oplus_{0\leqslant n\leqslant m}[G_{1}]_{n}\otimes_{\mathbb{k}}[G_{2}]_{m-n}$ for every $m\in{\mathbb{Z}}_{\geqslant 0}$. Thus, the images $[f]$, $[g]$ in $G$ are forms of degrees $a$ and $b$, respectively. By assumption one has that $[f][g]=[fg]$ vanishes, which is a contradiction. ∎
[…]
Let $R$ be a Noetherian domain and let $v$ be a discrete valuation of $R$. Let $m\in\mathbb{Z}_{>0}$ and consider the ideal $K=\{x\in R\mid v(x)\geqslant m\}$. Then $K$ is integrally closed.
Proof.
Assume that there exists $x\in\overline{K}\setminus K$, i.e., $0<v(x)<m$. Let $x^{n}+a_{1}x^{n-1}+\cdots+a_{n}=0$ with $a_{i}\in K^{i}$ be an integral equation of $x$. Since $(n-i)v(x)<(n-i)m$ for every $0\leqslant i<n$, we have $nv(x)<(n-i)m+iv(x)$. Therefore $v(x^{n}+a_{1}x^{n-1}+\cdots+a_{n})=nv(x)\in\mathbb{Z}_{>0}$, which is a contradiction. ∎
The following is the main ingredient in the proof of the main result of this section.
[…]
Adopt Section missing 4. Then $\left(\mathcal{B}\otimes\mathcal{C},(\underline{v},\underline{\gamma}),{\operatorname{conv}}(\Gamma,\Sigma)\right)$ is a Rees package for $IT+JT$.
Proof.
Set $\Omega:={\operatorname{conv}}(\Gamma,\Sigma)$. By Section missing 4 and Section missing 4 the $\mathcal{B}\otimes\mathcal{C}$-monomial functions $V_{k}$ in Section missing 3 corresponding to the non-coordinate supporting hyperplanes of $\Omega$ are valuations. Thus it suffices to show that for every $n\in\mathbb{Z}_{>0}$ the ideal $\overline{\left(IT+JT\right)^{n}}$ is spanned as a $\mathbb{k}$-vector space by $\mathcal{W}:=\{\mathfrak{b}\otimes\mathfrak{c}\in\mathcal{B}\otimes\mathcal{C}\mid(\underline{v}(\mathfrak{b}),\underline{\gamma}(\mathfrak{c}))\in n\Omega\}$. By the above conclusion on the non-coordinate supporting hyperplanes of $\Omega$ and by Section missing 4 we have that $\operatorname{span}(\mathcal{W})$ is an integrally closed ideal of $T$. Moreover, $\left(IT+JT\right)^{n}\subseteq\operatorname{span}(\mathcal{W})$ which follows from the fact that for every $w\in\mathbb{Q}_{>0}$ we have
(2) $w\Omega=\bigcup_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{R}}}}}\left(\alpha\Gamma+(w-\alpha)\Sigma\right).$
It remains to show $\operatorname{span}(\mathcal{W})\subseteq\overline{\left(IT+JT\right)^{n}}$. Let $\mathfrak{b}\otimes\mathfrak{c}\in\mathcal{W}$ so that $\underline{v}(\mathfrak{b})\in\alpha\Gamma$ and $\underline{\gamma}(\mathfrak{c})\in(n-\alpha)\Sigma$ for some $0\leqslant\alpha\leqslant n$, see (missing) 2. Let $d\in\mathbb{Z}$ be such that $\alpha d\in{\mathbb{Z}}_{\geqslant 0}$ and $(n-\alpha)d\in{\mathbb{Z}}_{\geqslant 0}$, then $\mathfrak{b}^{d}\in\overline{I^{\alpha d}}$ and $\mathfrak{c}^{d}\in\overline{J^{(n-\alpha)d}}$. Therefore $(\mathfrak{b}\otimes\mathfrak{c})^{d}\in\overline{I^{\alpha d}}\,\overline{J^{(n-\alpha)d}}\,T\subseteq\overline{\left(IT+JT\right)^{nd}}$, where the last inclusion follows from Section missing 2. We conclude that $\mathfrak{b}\otimes\mathfrak{c}\in\overline{\left(IT+JT\right)^{n}}$, finishing the proof. ∎
The following is the main result in this section. We note that, by Section missing 2, the sum on the right hand side is finite.
Theorem 4.1.
Adopt Section missing 4. For every $w\in\mathbb{Q}_{\geqslant 0}$ we have
(3) $\overline{\left(IT+JT\right)^{w}}=\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}\,T=\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{(IT)^{\alpha}}\,\overline{(JT)^{w-\alpha}}.$
Proof.
The first equality follows by combining Section missing 4, Section missing 3(2), and Equation (missing) 2. The second equality follows from Section missing 2. ∎
In the following corollary we highlight important classes of ideals for which the previous theorem applies.
[…]
Adopt Section missing 4. If $I$ and $J$ belong to any of the following classes of ideals, and not necessarily both from the same class, then (missing) 3 holds.
  1. (a)
    Monomial ideals in affine semigroup rings.
  2. (b)
    Sums of products of determinantal ideals of a generic matrix $X_{m\times n}$ and, if $\mathbb{k}=\mathbb{C}$, $\operatorname{GL}_{m}(\mathbb{C})\times\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[X_{m\times n}]$.
  3. (c)
    Sums of products of determinantal ideals of a generic symmetric matrix $Y_{n\times n}$ and, if $\mathbb{k}=\mathbb{C}$, $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[Y_{n\times n}]$.
  4. (d)
    Sums of products of ideals of Pfaffians of a skew-symmetric matrix $Z_{n\times n}$ and, if $\mathbb{k}=\mathbb{C}$, $\operatorname{GL}_{n}(\mathbb{C})$-invariant ideals of $\mathbb{C}[Z_{n\times n}]$.
  5. (e)
    Products of determinantal ideals of a Hankel matrix of variables.
Proof.
Combine Theorem 4.1 with: Theorem 3.1 for (a), Theorem 3.2, §missing 3.2 for (b), Theorem 3.3, §missing 3.3 for (c), Theorem 3.4, §missing 3.4 for (d), and Theorem 3.5 for (e). ∎
Next we see Theorem 4.1 and Section missing 4 in action.
[…]
Let $T=\mathbb{k}[x_{1}^{2}x_{2},x_{1}x_{2}^{3},y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}]$ and
$K=\left(x_{1}^{4}x_{2}^{2},x_{1}^{3}x_{2}^{4}\right)+\left(y_{1}y_{5}-y_{2}y_{4},y_{1}y_{6}-y_{3}y_{4},y_{2}y_{6}-y_{3}y_{5}\right)+\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)^{3}\subset T.$
Notice that $K=IT+JT$ where $I$ is the ideal in §missing 3.1 and $J$ the ideal $I_{\Lambda}$ in §missing 3.2. We identify these ideals with their images in $T$. We also set $J_{1}:=\left(y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}\right)$ and $J_{2}:=\left(y_{1}y_{5}-y_{2}y_{4},y_{1}y_{6}-y_{3}y_{4},y_{2}y_{6}-y_{3}y_{5}\right)$. By Section missing 4 we have
$\overline{K^{3/2}}=\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant 3/2,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{3/2-\alpha}}\,=\overline{I^{3/2}}+\overline{I}\overline{J^{1/2}}+\overline{I^{1/2}}\overline{J}+\overline{J^{3/2}},$
where the last equality is easily observed by inspecting Figure 1. Finally, using §missing 3.1, §missing 3.2, and Macaulay2 [29] we obtain
$\overline{K^{3/2}}$ $=(x_{1}^{6}x_{2}^{3},x_{1}^{5}x_{2}^{5})+(x_{1}^{4}x_{2}^{2},x_{1}^{3}x_{2}^{4})\overline{J^{1/2}}+(x_{1}^{2}x_{2})\overline{J}+\overline{J^{3/2}}$
$=(x_{1}^{6}x_{2}^{3},x_{1}^{5}x_{2}^{5})+(x_{1}^{4}x_{2}^{2},x_{1}^{3}x_{2}^{4})J_{1}^{2}+(x_{1}^{2}x_{2})\left(J_{1}^{3}+J_{2}\right)+\left(J_{2}^{(2)}+J_{1}^{2}J_{2}+J_{1}^{5}\right).$
Now we show that the conclusion of Theorem 4.1 does not always hold if the ideals $I$ and $J$ belong to the same $\mathbb{k}$-algebra.
[…]
Let $R=\mathbb{k}[x,y]$ be a polynomial ring over a field $\mathbb{k}$. Let $I=(xy^{3})$ and $J=(x^{3}y)$ respectively. For all $n\in\mathbb{Z}_{>0}$ we have
$x^{4n+2}y^{4n+2}\in\overline{\left(I\ +\,J\right)^{2n+1}}\setminus\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant 2n+1,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{2n+1-\alpha}}.$
From our methods in this section we also obtain the following interesting result. We recall the definition of the sum of two valuations from Section missing 4.
[…]
Adopt Section missing 4. If $v\in{\operatorname{\mathscr{RV}}}(I)$ and $\gamma\in{\operatorname{\mathscr{RV}}}(J)$ then $\gamma(J)v+v(I)\gamma\in{\operatorname{\mathscr{RV}}}(IT+JT)$. Conversely, every Rees valuation of $IT+JT$ can be obtained uniquely this way. In particular, $\left|{\operatorname{\mathscr{RV}}}(IT+JT)\right|=\left|{\operatorname{\mathscr{RV}}}(I)\right|\left|{\operatorname{\mathscr{RV}}}(J)\right|$.
Proof.
The result follows from Section missing 3(1), Section missing 4, and Section missing 4. ∎
[…]
Let $I=(x^{4})\subset\mathbb{k}[x]$, and let $J\subset\mathbb{k}[y_{1},y_{2},y_{3},y_{4},y_{5},y_{6}]$ be as in Section missing 4. By §missing 3.2, the polyhedra in the Rees packages of $I$ and $J$ are $\Gamma_{1}$ and $\Gamma_{2}$ as in Section missing 4. Thus by Section missing 4 (or Section missing 3) and Section missing 4, the Rees valuations of $IT+JT$ are the monomial functions $3\log_{x}+4\gamma_{1}+4\gamma_{2}$ and $\log_{x}+2\gamma_{1}$, where $\gamma_{1}$ and $\gamma_{2}$ are as in §missing 3.2, and $\log_{x}$ extracts the exponent of the variable $x$.
The previous corollary naturally gives rise to the following question.
[…]
Let $\mathbb{k}$ be an algebraically closed field. Let $R$ and $S$ be finitely generated $\mathbb{k}$-algebras that are domains and $T:=R\otimes_{\mathbb{k}}S$. Let $I\subset R$ and $J\subset S$ be nonzero ideals, do we always have $\left|{\operatorname{\mathscr{RV}}}(IT+JT)\right|=\left|{\operatorname{\mathscr{RV}}}(I)\right|\left|{\operatorname{\mathscr{RV}}}(J)\right|$?
In the next example we show that the answer to Section missing 4 is ‘no’ if we do not assume $\mathbb{k}=\overline{\mathbb{k}}$.
[…]
Consider the polynomial rings $R=\mathbb{Q}[x]$ and $S=\mathbb{Q}[y]$, and the prime ideals $I=(x^{2}+1)\subset R$, $J=(y^{2}+1)\subset S$. Since $\overline{R}\left[\dfrac{x^{2}+1}{x^{2}+1}\right]=\overline{R}=R$ and $IR$ has only one minimal prime we obtain $\left|{\operatorname{\mathscr{RV}}}(I)\right|=1$. Similarly $\left|{\operatorname{\mathscr{RV}}}(J)\right|=1$. Using Macaulay2 [29] we conclude that the ideal $IT+JT=(x^{2}+1,y^{2}+1)$ is integrally closed and that it has two minimal primes. Therefore $|\mathcal{R}\mathcal{V}(I+J)|\geqslant 2$ by [38, Discussion 10.1.3].
5. Asymptotic rational powers over algebraically closed fields
This section is devoted to the proof of Theorem 5.1 (or Theorem C). We will use the following notation throughout.
[…]
Let $\mathbb{k}$ be an algebraically closed field. Let $R$ and $S$ be $\mathbb{k}$-algebras, and $I\subset R$ and $J\subset S$ nonzero ideals. We denote by $IT+JT$ the $T$-ideal $IT+JT$, where $T:=R\otimes_{\mathbb{k}}S$.
The following is a main theorem of this section. We note that, by Section missing 2, the sum on the left hand side is finite.
Theorem 5.1.
Adopt Section missing 5 with $R$ and $S$ being finitely generated $\mathbb{k}$-algebras and normal domains. There exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}\,T\subseteq\overline{\left(IT+JT\right)^{w}}\subseteq\overline{I^{\tau}}T+\overline{J^{w-\tau}}T.$
[…]
We note that by [49, proof of Lemma 1.2] we have
$\sum_{\stackrel{{\scriptstyle 0\leqslant\tau\leqslant w,}}{{{\scriptscriptstyle\tau\in\mathbb{Q}}}}}\overline{I^{\tau}}\,\overline{J^{w-\tau}}\,T=\bigcap_{\stackrel{{\scriptstyle 0\leqslant\tau\leqslant w,}}{{{\scriptscriptstyle\tau\in\mathbb{Q}}}}}\left(\overline{I^{\tau}}T+\overline{J^{w-\tau}}T\right),$
[…]
Before presenting the proof of the theorem, we need some preparatory results. We start with the following statement, which is a generalization of [37, Theorem 2.3] that is suitable for our purposes. First, we need the following definitions.
[…]
Let $R$ be a regular domain and $\mathbf{x}=x_{1},\ldots,x_{d}$ a permutable regular sequence of $R$. We say $\mathbf{x}$ is a generalized regular system of parameters (grsop), if for every subset $\mathbf{x}^{\prime}=x_{i_{1}},\ldots,x_{i_{r}}$ of $\mathbf{x}$ the ring $R/(\mathbf{x}^{\prime})$ is also a regular domain. Let $\mathbf{x}$ be an grsop and let $I$ be an $R$-ideal generated by monomials in $\mathbf{x}$, i.e., $I=\left(\mathbf{x}^{\mathbf{a}_{1}},\ldots,\mathbf{x}^{\mathbf{a}_{s}}\right)$, where if $\mathbf{a}=(a_{1},\ldots,a_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$ then $\mathbf{x}^{\mathbf{a}}:=x_{1}^{a_{1}}\cdots x_{d}^{a_{d}}$. The Newton polyhedron of $I$ is defined as
${\operatorname{NP}}(I):={\operatorname{conv}}\left(\mathbf{a}_{1},\ldots,\mathbf{a}_{s}\right)+\mathbb{R}^{d}_{\geqslant 0}\subseteq\mathbb{R}^{d}.$
Theorem 5.2.
Let $R$, $\mathbf{x}$, and $I$ be as in Section missing 5. For every $w\in\mathbb{Q}_{\geqslant 0}$ the ideal $\overline{I^{w}}$ is generated by the monomials $\mathbf{x}^{\mathbf{e}}$ with $\mathbf{e}\in w{\operatorname{NP}}(I)$.
Proof.
We only include relevant details of how a modification of the proof of [37, Theorem 2.3] leads to our statement. We refer the reader to the proof of [37, Theorem 2.3] for any undefined notation, which we will follow closely.
Let $\mathbf{e}\in w{\operatorname{NP}}(I)$ and $c_{1},\ldots,c_{s}\in\mathbb{Q}_{\geqslant 0}$ be such that $\sum_{i}c_{i}=w$ and $\mathbf{e}\geqslant\sum_{i}c_{i}\mathbf{a}_{i}$ componentwise. Write $c_{i}=\frac{m_{i}}{n}$ and $w=\frac{p}{q}$ where $m_{i},p\in{\mathbb{Z}}_{\geqslant 0}$ and $n,q\in\mathbb{Z}_{>0}$. Therefore, $qn\mathbf{e}\geqslant\sum_{i}qm_{i}\mathbf{a}_{i}$. It follows that
$\left(\mathbf{x}^{q\mathbf{e}}\right)^{n}\in I^{\sum_{i}qm_{i}}=\left(I^{p}\right)^{n},$
and then $\left(\mathbf{x}^{\mathbf{e}}\right)^{q}\in\overline{I^{p}}$, i.e., $\mathbf{x}^{\mathbf{e}}\in\overline{I^{w}}$.
We proceed with the reverse inclusion. Let $H$ be a non-coordinate supporting hyperplane of ${\operatorname{NP}}(I)$ defined by $\langle\mathbf{h},\mathbf{X}\rangle=h$ where $\mathbf{h}=(h_{1},\ldots,h_{d})\in{\mathbb{Z}}_{\geqslant 0}^{d}$, $h\in\mathbb{Z}_{>0}$, and $\mathbf{X}=(X_{1},\ldots,X_{d})$. Let $I_{H}$ be the ideal generated by the monomials $\mathbf{x}^{\mathbf{e}}$ with $\langle\mathbf{h},\mathbf{e}\rangle\geqslant h$. It suffices to show that the conclusion of the theorem holds for $I_{H}$. Let $\mathfrak{p}=(Y_{1},\ldots,Y_{t})R^{\prime}$. By Section missing 2 and [38, Corollary 5.4.2], the rational powers of $\mathfrak{p}^{h}R^{\prime}_{\mathfrak{q}}$ have no embedded primes for every prime ideal $\mathfrak{p}\subseteq\mathfrak{q}$. Thus, it suffices to show that $\overline{\left(\mathfrak{p}^{h}R^{\prime}_{\mathfrak{p}}\right)^{w}}=\mathfrak{p}^{\lceil wh\rceil}R^{\prime}_{\mathfrak{p}}$, which follows from $\mathfrak{p}R^{\prime}_{\mathfrak{p}}$ being a prime ideal generated by a regular sequence, as in this case the powers $\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}$ are integrally closed and $\operatorname{gr}_{R^{\prime}_{\mathfrak{p}}}(\mathfrak{p}R^{\prime}_{\mathfrak{p}}):=\oplus_{n\in{\mathbb{Z}}_{\geqslant 0}}\mathfrak{p}^{n}R^{\prime}_{\mathfrak{p}}/\mathfrak{p}^{n+1}R^{\prime}_{\mathfrak{p}}$ is a domain. ∎
[…]
Let $\mathbb{k}$ be an algebraically closed field. Let $(V_{1},\eta_{1})$ and $(V_{2},\eta_{2})$ be two divisorial valuation rings containing $\mathbb{k}$ and such that $V:=V_{1}\otimes_{\mathbb{k}}V_{2}$ is Noetherian. Let $I=(\eta_{1}^{n_{1}})$ and $J=(\eta_{2}^{n_{2}})$ be ideals of $V_{1}$ and $V_{2}$, respectively. For every $\tau,w\in\mathbb{Q}_{\geqslant 0}$ with $\tau\leqslant w$, we have the following inclusion of $V$-ideals
$\overline{\left(IT+JT\right)^{w}}\subseteq\overline{I^{\tau}}T+\overline{J^{w-\tau}}T.$
Proof.
By [57, Theorem 6(e)] and [5, Chapter V, §17.2, Corollary], $V$ is a regular domain and $\eta_{1},\eta_{2}$ is a grsop of $V$ (see Section missing 5). Thus by Theorem 5.2 the ideal $\overline{\left(IT+JT\right)^{w}}$ is generated by the monomials $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ with $n_{2}e_{1}+n_{1}e_{2}\geqslant wn_{1}n_{2}$. Therefore, for any such $\eta_{1}^{e_{1}}\eta_{2}^{e_{2}}$ we must have $e_{1}\geqslant\tau n_{1}$ or $e_{2}\geqslant(w-\tau)n_{2}$. The result follows as $\overline{I^{\tau}}$ is generated by $\eta_{1}^{\lceil\tau n_{1}\rceil}$ and $\overline{J^{w-\tau}}$ by $\eta_{2}^{\lceil(w-\tau)n_{2}\rceil}$. ∎
The following lemma offers a vanishing result for rational powers of exceptional divisors of normalized blowups that can be of independent utility and interest.
[…]
Adopt Section missing 2. Let $v:X^{+}\longrightarrow X$ be the normalization of the blowup of $X$ along $\mathcal{I}$, and let $E$ be its exceptional divisor, i.e., $\mathcal{I}\cdot\mathcal{O}_{X^{+}}=\mathcal{O}_{X^{+}}(-E)$. Then there exists a positive integer $n$ such that if $\beta\in\mathbb{Q}$ and $\beta\geqslant n$ we have $R^{1}v_{*}\left(\mathcal{O}_{X^{+}}(-\lceil\beta E\rceil)\right)=0$.
Proof.
Set $\mathscr{T}:=\mathcal{O}_{X^{+}}(-E)$. By Section missing 2 we have $\overline{\mathscr{T}^{\beta}}=\mathcal{O}_{X^{+}}\left(-\lceil\beta E\rceil\right)$ for every $\beta\in\mathbb{Q}_{\geqslant 0}$. Notice that there exists $m\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant m$ we have $\overline{\left(\mathscr{T}^{+}\right)^{\beta+m}}=\overline{\left(\mathscr{T}^{+}\right)^{\beta}}\,\,\overline{\left(\mathscr{T}^{+}\right)^{m}}$. Indeed, for any finite affine cover $\{U_{\lambda}\}_{\lambda\in\Lambda}$ of $X^{+}$, this holds locally on each $U_{\lambda}$ by Section missing 2(3), producing finitely many $m_{\lambda}$. So, we take can take $m:={\operatorname{lcm}}(m_{\lambda}\mid\lambda\in\Lambda)$. Therefore for every $\beta\geqslant m$ we have
(4) $\overline{\left(\mathscr{T}^{+}\right)^{\beta}}=\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\frac{\beta}{m}\rfloor+m}}\,\,\left(\overline{\left(\mathscr{T}^{+}\right)^{m}}\right)^{\lfloor\frac{\beta}{m}\rfloor-1}\cong\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\beta/m\rfloor+m}}\otimes\mathcal{O}_{X^{+}}(-E)^{\otimes m\lfloor\beta/m\rfloor-m},$
where the last isomorphism holds by flatness. Since $m\leqslant\beta-m\lfloor\beta/m\rfloor+m\leqslant 2m$, by Section missing 2 there are finitely many distinct ideal sheaves of the form $\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\beta/m\rfloor+m}}$. Since $\mathcal{O}_{X^{+}}(-E)$ is ample, by Serre’s vanishing theorem [33, Proposition III.5.3] there exists $n\in\mathbb{Z}_{>0}$ such that for every $\beta\geqslant n$
(5) $R^{1}v_{*}\left(\overline{\left(\mathscr{T}^{+}\right)^{\beta-m\lfloor\beta/m\rfloor+m}}\otimes\mathcal{O}_{X^{+}}(-E)^{\otimes m\lfloor\beta/m\rfloor-m}\right)=0.$
The result now follows by combining (missing) 4 and (missing) 5. ∎
We are now ready to prove the main theorem of this section.
Proof of Theorem 5.1.
Fix $\alpha,w\in\mathbb{Q}_{\geqslant 0}$ with $0\leqslant\alpha\leqslant w$. By Section missing 2 we have
$\overline{I^{\alpha}}\,\overline{J^{w-\alpha}}T=\overline{(IT)^{\alpha}}\,\,\overline{(JT)^{w-\alpha}}\subseteq\overline{(IT+JT)^{w}},$
proving the first inclusion. We proceed with the second inclusion to show that there exists $w_{0}\in\mathbb{Q}_{>0}$ such that
(6) $\overline{\left(IT+JT\right)^{w}}\subseteq\overline{I^{\tau}}T+\overline{J^{w-\tau}}T$
for all $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$. Let $X_{1}={\operatorname{Spec\,}}(R)$ and $X_{2}={\operatorname{Spec\,}}(S)$ so that $X_{1}\times X_{2}={\operatorname{Spec\,}}(T)$. Let $\mathcal{I}=\tilde{I}\subseteq O_{X_{1}}$ and $\mathcal{T}=\tilde{J}\subseteq O_{X_{2}}$ be the sheaves associated to $I$ and $J$, respectively. Let $v_{1}:X_{1}^{+}\longrightarrow X_{1}$ and $v_{2}:X_{2}^{+}\longrightarrow X_{2}$ be the normalization of the blowups of $X_{1}$ and $X_{2}$ along $\mathcal{I}$ and $\mathcal{T}$, respectively, so that $\mathcal{I}^{+}:=\mathcal{I}\cdot\mathcal{O}_{X_{1}^{+}}=:\mathcal{O}_{X_{1}^{+}}(-E_{1})$ and $\mathcal{T}^{+}:=\mathcal{T}\cdot\mathcal{O}_{X_{2}^{+}}=:\mathcal{O}_{X_{2}^{+}}(-E_{2})$. Moreover, consider $Y:=X_{1}\times X_{2}$, $Y^{+}:=X_{1}^{+}\times X_{2}^{+}$, the canonical projections
$p_{1}:Y\longrightarrow X_{1},\,\,\,\,p_{2}:Y\longrightarrow X_{2},\,\,\,\,q_{1}:Y^{+}\longrightarrow X_{1}^{+},\,\,\,\,q_{2}:Y^{+}\longrightarrow X_{2}^{+},$
and $f:=v_{1}\times v_{2}:Y^{+}\longrightarrow Y$. These maps are represented in the following diagram.
[…]
We note that it suffices to show that the following two claims are true. Here, for simplicity of notation, for an ideal sheaf $\mathcal{J}\subseteq O_{X_{i}^{+}}$, resp. $\mathcal{J}\subseteq O_{X_{i}}$, we write $q_{i}^{-1}\mathcal{J}$, resp. $p^{-1}_{i}\mathcal{J}$, for the ideal sheaf $q_{i}^{-1}\mathcal{J}\cdot\mathcal{O}_{Y^{+}}$, resp. $p_{i}^{-1}\mathcal{J}\cdot\mathcal{O}_{Y}$. We note that, since each $q_{i}$, $p_{i}$ is flat, the latter are isomorphic to $q_{i}^{*}\mathcal{J}$, resp. $p_{i}^{*}\mathcal{J}$.
Claim 1: For any rational number $0\leqslant\tau\leqslant w$ we have
$\overline{\left(q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}\right)^{w}}\subseteq q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}+q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}.$
Claim 2: There exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$f_{*}\left(q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}+q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\right)=f_{*}\left(q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\right)+f_{*}\left(q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\right).$
Indeed, if both claims hold, since $q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}=f^{-1}\left(p_{1}^{-1}\mathcal{I}+p_{2}^{-1}\mathcal{T}\right)\cdot\mathcal{O}_{Y^{+}}$, then from Theorem 2.2 one obtains
$\overline{(p_{1}^{-1}\mathcal{I}+p_{2}^{-1}\mathcal{T})^{w}}=f_{*}\left(\overline{\left(q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}\right)^{w}}\right)$ $\subseteq f_{*}\left(q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\right)+f_{*}\left(q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\right)$
$=p_{1}^{-1}\overline{\left(\mathcal{I}\right)^{\tau}}+p_{2}^{-1}\overline{\left(\mathcal{T}\right)^{w-\tau}},$
for $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$. The last equality follows from Künneth formula (see [55, tag 0BEC, Lemma 33.29.1] and [34, Proposition III.8.5]) using Theorem 2.2 and the facts that $v_{1*}\left(\mathcal{O}_{X_{1}^{+}}\right)=\mathcal{O}_{X_{1}}$ and $v_{2*}\left(\mathcal{O}_{X_{2}^{+}}\right)=\mathcal{O}_{X_{2}}$ [34, proof of Corollary III.11.4]. Thus, (missing) 6 will follow by taking global sections, see Section missing 2.
We proceed to prove Claim 1, for which we can restrict to a product of affine schemes $U_{1}\times U_{2}\subset Y^{+}$, where $U_{1}={\operatorname{Spec\,}}(A_{1})\subset X_{1}^{+}$, $V={\operatorname{Spec\,}}(A_{2})\subset X_{2}^{+}$, and $A_{1}$ and $A_{2}$ are Noetherian normal domains. Therefore $\mathcal{I}^{+}$ and $\mathcal{T}^{+}$ can be identified with principal ideals $(a_{1})\subseteq A_{1}$ and $(a_{2})\subseteq A_{2}$, respectively, and $q_{1}^{-1}\mathcal{I}^{+}+q_{2}^{-1}\mathcal{T}^{+}$ with $(a_{1})A+(a_{2})A$ where $A:=A_{1}\otimes_{\mathbb{k}}A_{2}$. By Section missing 2 and [38, Corollary 5.4.2] the associated primes of the rational powers $\overline{(a_{1})^{\tau}}$ and $\overline{(a_{2})^{w-\tau}}$ all have codimension one. By [32, Corollary 2.8] the associated primes of $(a_{1})A+(a_{2})A$ are sums $\mathfrak{p}_{1}A+\mathfrak{p}_{2}A$ with $\mathfrak{p}_{i}$ associated prime of $(a_{i})$, $i=1,2$. Moreover, rational powers localize by Section missing 2. Thus, to show $\overline{\left((a_{1})A+(a_{2})A\right)^{w}}\subseteq\overline{(a_{1})^{\tau}}A+\overline{(a_{2})^{w-\tau}}A$, it suffices to do so in the localizations $(A_{1})_{\mathfrak{p}_{1}}\otimes_{\mathbb{k}}(A_{2})_{\mathfrak{p}_{2}}$. But the latter follows from Section missing 5, finishing the proof of Claim 1.
We continue with the proof of Claim 2. Consider the exact sequence
$0\longrightarrow q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\cap q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}$ $\longrightarrow q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\oplus q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}$
$\longrightarrow q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}+q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\longrightarrow 0.$
Applying $f_{*}$, the claim reduces to showing
$R^{1}f_{*}\left(q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\cap q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\right)=R^{1}f_{*}\left(q_{1}^{-1}\overline{\left(\mathcal{I}^{+}\right)^{\tau}}\otimes q_{2}^{-1}\overline{\left(\mathcal{T}^{+}\right)^{w-\tau}}\right)=0.$
By Künneth formula it suffices to show that there exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $\beta\in\mathbb{Q}$ and $\beta\geqslant w_{0}$ then
(7) $R^{1}v_{1*}\left(\overline{\left(\mathcal{I}^{+}\right)^{\beta}}\right)=0\qquad\text{and}\qquad R^{1}v_{2*}\left(\overline{\left(\mathcal{T}^{+}\right)^{\beta}}\right)=0.$
The latter vanishings hold by Section missing 5, concluding the proof of Claim 2. This finishes the proof of the theorem. ∎
[…]
We note that if for certain ideals $I$ and $J$ the vanishings in (7) hold for every $\beta\geqslant 0$, then the proof of Theorem 5.1 shows that these ideals satisfy the conclusion of Section missing 1.
By working locally in affine open subsets, Theorem 5.1 implies the following more general statement (cf. [49, Theorem 0.3]). As in the proof of Theorem 5.1, we denote by $p^{-1}\mathcal{J}$, resp. $q^{-1}\mathcal{J}$, the ideal sheaves $p^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$, resp. $q^{-1}\mathcal{J}\cdot\mathcal{O}_{X\times Y}$.
[…]
Let $X,Y$ be normal varieties over an algebraically closed field $\mathbb{k}$. Let $\mathcal{I}\subseteq\mathcal{O}_{X},\mathcal{T}\subseteq\mathcal{O}_{y}$ be nonzero ideal sheaves, and $p:X\times Y\longrightarrow X$, $q:X\times Y\longrightarrow Y$ the canonical projections. There exists $w_{0}\in\mathbb{Q}_{>0}$ such that if $w_{0}\leqslant\tau\leqslant w-w_{0}$, $\tau\in\mathbb{Q}$ then
$\sum_{\stackrel{{\scriptstyle 0\leqslant\alpha\leqslant w,}}{{{\scriptscriptstyle\alpha\in\mathbb{Q}}}}}p^{-1}\overline{\mathcal{I}^{\alpha}}\,\,q^{-1}\overline{\mathcal{T}^{w-\alpha}}\subseteq\overline{\left(p^{-1}\mathcal{I}\,+\,q^{-1}\mathcal{T}\right)^{w}}\subseteq p^{-1}\overline{\left(\mathcal{I}\right)^{\tau}}\,+\,q^{-1}\overline{\left(\mathcal{T}\right)^{w-\tau}}.$
Similarly as in [49, Proposition 2.1], the formula in Theorem 5.1 can be used to approximate rational powers up to $\mathfrak{m}$-adic neighborhoods. This is also related to the main theorem in [23, 24].
[…]
Adopt Section missing 5 with $R$ a finitely generated $\mathbb{k}$-algebra and normal domain, and $S$ a polynomial ring with homogeneous maximal ideal $\mathfrak{m}$. For every $n\in{\mathbb{Z}}_{\geqslant 0}$, there exists $w_{0}=w_{0}(n)\in\mathbb{Q}_{>0}$ such that if $w,\ell\geqslant w_{0}$ then
$\overline{\left(IT+\mathfrak{m}^{n}T\right)^{w+\ell}}\subseteq\overline{I^{w}}T+\mathfrak{m}^{\lceil n\ell\rceil}T.$
Moreover, for any nonzero ideal $J\subset S$ there exists a $w_{0}\in\mathbb{Q}_{>0}$ such that if $w,\ell\geqslant w_{0}$ then
$\overline{\left(IT+JT\right)^{w+\ell}}\subseteq\overline{I^{w}}T+J^{\lfloor\ell\rfloor-\dim(S)+1}T.$
Proof.
The first statement follows from Theorem 5.1 and since the powers of $\mathfrak{m}$ are integrally closed, then $\overline{\left(\mathfrak{m}^{n}\right)^{\ell}}=\mathfrak{m}^{\lceil n\ell\rceil}$. The second statement also follows from Theorem 5.1 and by the inclusions $\overline{J^{\ell}}\subseteq\overline{J^{\lfloor\ell\rfloor}}\subseteq J^{\lfloor\ell\rfloor-\dim(S)+1}$, which hold by the Briançon-Skoda Theorem [47] (see also [38, Corollary 13.3.4]). ∎